### v1

# Masked Generative Extractor for Synergistic Representation and 3D Generation of Point Clouds

Abstract
In the field of 2D image generation modeling and representation learning, Masked Generative Encoder (MAGE) has demonstrated the synergistic potential between generative modeling and representation learning. Inspired by this, we propose Point-MAGE to extend this concept to point cloud data. Specifically, this framework first utilizes a Vector Quantized Variational Autoencoder (VQVAE) to reconstruct a neural field representation of 3D shapes, thereby learning discrete semantic features of point patches. Subsequently, by combining the masking model with variable masking ratios, we achieve synchronous training for both generation and representation learning. Furthermore, our framework seamlessly integrates with existing point cloud self-supervised learning (SSL) models, thereby enhancing their performance. We extensively evaluate the representation learning and generation capabilities of Point-MAGE. In shape classification tasks, Point-MAGE achieved an accuracy of 94.2% on the ModelNet40 dataset and 92.9% (+1.3%) on the ScanObjectNN dataset. Additionally, it achieved new state-of-the-art performance in few-shot learning and part segmentation tasks. Experimental results also confirmed that Point-MAGE can generate detailed and high-quality 3D shapes in both unconditional and conditional settings.
1 Introduction
Recent studies have confirmed that representation learning frameworks based on masked modeling Yu_2022_CVPR ; liu2022masked ; pang2022masked ; zhang2022point ; NEURIPS2023_5ed5c3c8 can extract high-quality 3D representations. At the same time, generative models zhang20223dilg ; Mittal_2022_CVPR ; Yan_2022_CVPR ; zhang20233dshape2vecset have also made significant achievements in generating realistic 3D shapes. However, there is little research that attempts to combine the potential of these two types of models to explore their joint application in 3D shapes. We note that in the field of 2D images, Masked Generative Encoder (MAGE) Li_2023_CVPR proposed an innovative unified framework for generation training and representation learning, and experimentally verified the synergistic effects between these two tasks. This inspired us to pose a question: Can the characteristics demonstrated in the MAGE framework be extended to point cloud data to achieve deeper 3D representation learning and generation capabilities?
The strategy of MAGE Li_2023_CVPR is to combine Masked Image Modeling (MIM) with variable masking ratios. Specifically, this method conducts generation training under high masking ratios, while representation learning is performed under low masking ratios. Following this idea, we explore combining masked point cloud modeling with variable masking ratios to realize the Point-MAGE framework. However, due to the fundamental differences in structure and distribution between images and point cloud data, this direct transfer faces several challenges.
Firstly, MAGE Li_2023_CVPR effectively obtains discrete semantic features from images through pixel-wise reconstruction training. However, directly using point clouds as reconstruction targets introduces sampling bias. This is because 3D shapes are inherently continuous, while point clouds obtained through sampling merely represent them discretely. This leads to non-zero loss between different point cloud samples of the same object, affecting the accurate understanding of shapes. Moreover, methods solely relying on point cloud reconstruction fail to acquire more advanced representations of 3D shapes, such as neural radiance fields or triangle meshes, thereby limiting the potential of models in generation tasks. Therefore, simply applying variable masking ratios to existing point cloud SSL methods Yu_2022_CVPR ; pang2022masked ; zhang2022point ; NEURIPS2023_5ed5c3c8 cannot achieve the goals of Point-MAGE, as these existing methods typically use Chamfer distance (CD) as the reconstruction loss for point cloud reconstruction. To address these challenges, we instead focus the reconstruction targets on 3D shapes represented as neural radiance fields (NeRF) and extract discrete semantic features of point clouds through a VQVAE van2017neural . Point-MAGE leverages these high-level semantic discrete features as input and reconstruction targets, learning the probability distribution of masked tokens. After training, our method not only iteratively performs three-dimensional shape generation tasks but also resolves sampling bias in representation learning through the vector quantization process. This is because different point cloud samples from the same local block are unified into the same discrete token during the vector quantization stage, ensuring consistency of representations.
Secondly, the inherent unordered nature of point cloud poses a challenge for Transformer-based generative models, as these models require serialized inputs. To address this issue, we partition the complete point cloud into multiple patches, each defined by a fixed number of points and a central point. To serialize these patches, we adopt the Morton sorting algorithm used in PointGPT NEURIPS2023_5ed5c3c8 , which can sort point cloud patches while preserving geometric relationships. In generation task, since the central point of point cloud patch is unknown, we introduce relative position encoding in the generator of Point-MAGE as a replacement for traditional absolute position encoding, and use the central points as the reconstruction target.
Finally, we conducted an analysis of the synergistic effects of the two tasks. We found that generative models benefit from the variable masking ratio. Unlike traditional methods that reconstruct all tokens from scratch, our approach leverages a small number of visible tokens to provide additional sample information, which not only accelerates convergence but also enhances the model’s final performance. For representation learning, by shifting the reconstruction target from raw point clouds to discrete semantic features, our framework successfully overcomes the variance issue in point cloud sampling and focuses more on the global features of 3D shapes. Furthermore, the Point-MAGE framework can be integrated with SSL model designs proposed in previous work, further enhancing the performance of representation learning. Experimental results demonstrate the effectiveness of this strategy, with Point-MAGE achieving state-of-the-art (SOTA) performance on multiple downstream tasks. Specifically, our method achieved an accuracy of 94.2% on the ModelNet40 dataset Wu_2015_CVPR and 92.9% on the ScanObjectNN dataset yi2016scalable , respectively, for shape classification tasks. In the part segmentation task on the ShapeNetPart yi2016scalable dataset, our method achieved a mean class Intersection over Union (IoU) of 85.0%. Additionally, we demonstrated the pre-trained model’s ability to generate realistic 3D shapes in both unconditional and conditional settings, further showcasing the flexibility and powerful generative capabilities of Point-MAGE.
Our main contributions can be summarized as follows: (I) We propose Point-MAGE, a unified framework for point cloud data representation learning and generative training that combines a two-stage training method with a variable masking ratio design. (II) Point-MAGE overcomes the variance issue in point cloud sampling and can be integrated with model architectures from previous SSL works, setting a new SOTA benchmark for downstream tasks such as shape classification, few-shot learning, and part segmentation. (III) Point-MAGE is capable of generating high-quality 3D shapes in both unconditional and conditional settings.
2 Related Works
Representation Learning for Point Clouds.
In recent years, self-supervised frameworks for point cloud representation learning have made significant progress, forming two main model paradigms. Firstly, methods based on contrastive learning Yang_2018_CVPR ; xie2020pointcontrast ; Navaneet_2020_CVPR ; Zhang_2021_ICCV ; Afham_2022_CVPR have played an important role in point cloud self-supervised learning (SSL). These methods can explore the intrinsic properties of point clouds, but they often rely on cross-modal information or teacher models, which undoubtedly add extra burden to the training process. Secondly, more closely related to our work is another class of methods based on masked point cloud modeling. Inspired by the successful experience of masked modeling in natural language processing (NLP) radford2018improving ; devlin2018bert ; radford2019language ; brown2020language and image processing he2022masked ; Wei_2022_CVPR ; zhou2021ibot fields, recent research has focused on randomly masking point cloud regions and reconstructing these regions Yu_2022_CVPR ; pang2022masked ; zhang2022point ; NEURIPS2023_5ed5c3c8 , prompting the encoder to learn high-quality point cloud representations. In particular, Point-M2AE zhang2022point proposed an innovative hierarchical transformer structure, extending Point-MAE pang2022masked to capture multi-scale features of point clouds. On the other hand, PointGPT attempts to solve the problem of global shape leakage through autoregressive generation, but it does not have the ability to generate high-quality 3D models from scratch. Additionally, methods based on masked region reconstruction are susceptible to point cloud sampling variance, which may introduce additional understanding and bias issues. To overcome these challenges, our method adopts a strategy of masked modeling in the high-level semantic space, effectively addressing the aforementioned issues.
Generative Models for 3D Shapes.
In the field of 3D generation, Generative Adversarial Networks (GANs) wu2016learning ; achlioptas2018learning ; Chen_2019_CVPR ; Ibing_2021_CVPR ; zheng2022sdf constitute a major generative paradigm, despite their limitations in training stability and mode collapse issues. Meanwhile, some research has turned to alternative approaches such as Normalizing Flows (NFs) dinh2014nice ; rezende2015variational ; yang2019pointflow , Energy-Based Models (EBMs) lecun2006tutorial ; xie2016theory ; Xie_2021_CVPR ; xie2020generative , and Autoregressive Models (ARs) Mittal_2022_CVPR ; cheng2022autoregressive ; Yan_2022_CVPR ; zhang20223dilg . Recently, inspired by the significant progress of Diffusion Models (DMs) dhariwal2021diffusion ; song2020score in image synthesis, some studies have begun to explore the application of diffusion models to 3D shape generation vahdat2022lion ; hui2022neural ; Shue_2023_CVPR ; Chou_2023_ICCV ; zhang20233dshape2vecset . Although these models can generate realistic 3D models, they often lack the ability to extract high-quality semantic representations from 3D shapes, limiting their direct application in point cloud representation learning tasks. To address this challenge, we propose an innovative approach aimed at generating diverse samples while extracting high-quality semantic embeddings from 3D shapes to promote data generalization and enhance the performance of downstream tasks. We draw on the successful experience of MAGE Li_2023_CVPR in image synthesis and representation learning, and realize the organic combination of generation and representation learning on point clouds through a two-stage scheme. In the first stage, we use VQVAE van2017neural to train an encoder and decoder to obtain latent codes. Subsequently, in the second stage, we further train the extractor for representation learning through masked modeling and use a bidirectional transformer generator to complete the generation task.
3 Method
Point-MAGE is designed as a unified framework for both point cloud generation tasks and representation learning. As shown in Figure 1, the framework first divides the input point cloud into multiple local point blocks and serializes them based on their geometric positional relationships. Then, it uses a mini-PointNet Qi_2017_CVPR model for initial feature extraction of each point block. Subsequently, by training a VQVAE van2017neural model, these point block features are quantized into discrete semantic latents. Finally, the framework combines a variable masking ratio with a Vision Transformer (ViT)-based extractor-generator architecture to predict masked semantic tokens.
Figure 1: Overall pipeline of Point-MAGE. The training process consists of two stages. In the first stage, a VQVAE van2017neural is used to reconstruct 3D shapes represented by NeRF, converting the input point cloud into a series of discrete semantic latents. The second stage is a masked token reconstruction process, where a ViT-based extractor-generator architecture is employed to extract high-quality feature representations from the unmasked tokens and to reconstruct the masked tokens.
3.1 Point Patch Embedding
Point Cloud Patching.
We determine the number of patches and the number of points per patch in the point cloud sampling process through two key hyperparameters $G$ and $K$. Specifically, given the input point cloud $X\in\mathbb{R}^{n\times 3}$, we use the farthest point sampling (FPS) strategy to select $G$ central points,
$C=\{x_{i}\}_{i\in\mathcal{N}}=\text{FPS}\left(X\right),~{}C\in\mathbb{R}^{G\times 3},$ (1)
where $|\mathcal{N}|=n$. Then, using the K-nearest neighbors (KNN) algorithm, we identify $|\mathcal{N}_{i}|=K$ neighbors of the central point $\{x_{i}\}_{i\in\mathcal{N}}$, resulting in $G$ point cloud patches,
$P=\{x_{j}\}_{j\in\mathcal{N}_{i}}=\text{KNN}\left(C,X\right),~{}P\in\mathbb{R}^{G\times K\times 3}.$ (2)
Embedding.
To provide serialized input, we follow NEURIPS2023_5ed5c3c8 and encode the coordinates of the central points $x_{i}$ into one-dimensional space using Morton ordering morton1966computer , and then sort them to obtain serialized central points $C^{s}\in\mathbb{R}^{G\times 3}$ and point cloud patches $P^{s}\in\mathbb{R}^{G\times 3}$. We project each patch into an embedding vector using the mini-PointNetQi_2017_CVPR module,
[Uncaptioned image]${\color[rgb]{0.875,0.4765625,0.3671875}~{}:P^{s}_{E}}=\{e_{i}\}_{i\in\mathcal{G}}=\text{PointNet}\left(P^{s}\right),~{}P^{s}_{E}\in\mathbb{R}^{G\times D},$ (3)
where $|\mathcal{G}|=G$, D is the embedding dimension of the point patch.
3.2 Tokenization
We train a VQVAE van2017neural to obtain a quantizer that embeds input point patches into a series of semantic latents. This tokenization step enables our model to focus on high-level semantics, which helps to overcome the variance issue in point cloud sampling and allows for a more comprehensive understanding of 3D shapes.
Vector Quantization.
We use a ViT-based encoder to learn the local latents $T$. The input to the transformer includes the sorted point block embeddings $P^{s}_{E}$ and the sinusoidal positional encodings vaswani2017attention $C^{s}_{E}$ of the center point,
$T=\{t_{i}\}_{i\in\mathcal{G}}=\text{Encoder}\left(P^{s}_{E},C^{s}_{E}\right),~{}T\in\mathbb{R}^{G\times D}.$ (4)
The output of the encoder is then replaced with the closest vector in the codebook $\mathcal{Q}$ through vector quantization,
$T^{Q}=\{t^{q}_{i}\}_{i\in\mathcal{G}}=\mathop{\arg\min}\limits_{t^{q}_{j}\in\mathcal{Q}}\|t^{q}_{j},t_{i}\|,~{}T^{Q}\in\mathbb{R}^{G\times D}.$ (5)
NeRF Reconstruction.
We utilize a ViT-based decoder along with a reconstruction head to reconstruct 3D shapes represented in NeRF. The decoder takes as input the quantized latents and the position encoding of the center points,
$T^{D}=\{t^{d}_{i}\}_{i\in\mathcal{G}}=\text{Decoder}\left(T^{Q},C^{s}_{E}\right),~{}T^{D}\in\mathbb{R}^{G\times D}.$ (6)
For any $x\in\mathbb{R}^{3}$, we use the Nadaraya-Watson estimator for interpolation to obtain its latent $t^{d}_{x}$. Then, we estimate $\mathcal{O}(x)$ through a multi-layer perceptron and a sigmoid activation function,
$\mathcal{O}(x)=\text{Sigmoid}\left(\text{MLP}\left(x,t^{d}_{x}\right)\right),~{}\mathcal{O}(x)\in\left[0,1\right].$ (7)
The optimization objective consists of binary cross-entropy (BCE) loss and an additional commitment loss,
$\mathcal{L}_{n}=\mathbb{E}_{x\in\mathbb{R}^{3}}[\text{BCE}(\mathcal{O}(x),\hat{\mathcal{O}}(x))]+\lambda\mathbb{E}_{x\in\mathbb{R}^{3}}[\mathbb{E}_{i\in\mathcal{M}}\|\epsilon(t^{q}_{i})-t_{i}\|^{2}],$ (8)
where $\hat{\mathcal{O}}(x)$ denotes the ground truth, and $\epsilon(\cdot)$ represents the stop-gradient operation. In this paper, $\lambda=0.25$ is a weighting coefficient.
3.3 Pre-training
After obtaining the tokenizer, we simultaneously train an extractor and a generator model through masking modeling. Below, we introduce the masking strategy, model design, and optimization objectives of Point-MAGE.
Masking Strategy.
For a given point cloud $X\in\mathbb{R}^{n\times 3}$, we first use the trained tokenizer to obtain the corresponding discrete tokens $T^{Q}$. Subsequently, we introduce a variable mask ratio to apply masking, which can be calculated as:
$m_{r}\sim N(\mu=0.55,\sigma^{2})~{}\text{with}~{}m_{r}\in\left[0.5,1\right]$ (9)
Where $N$ represents the truncated Gaussian distribution. After masking, to reduce training time and memory consumption, we randomly drop a portion of masked tokens, ensuring that only $0.5\times G$ tokens are input to the extractor, following the MAGE Li_2023_CVPR .
Extractor and Generator.
We feed the remaining tokens into the ViT-based extractor-generator structure. Then, we concatenate a learnable mask token to the input sequence. The extractor processes the input tokens $T^{Q}_{m}$ and encodes them into the latent feature space,
$T^{E}=\{t^{e}_{i}\}_{i\in\mathcal{G},m_{i}=0}=\text{Extractor}(T^{Q}_{m},C^{s}_{E}),~{}T^{E}\in\mathbb{R}^{\frac{G}{2}\times D},$ (10)
where $M=[m_{i}]^{G}_{i=1}$ represents a binary mask indicating whether a token is masked or not. Then, the output of the extractor is padded to the complete length $G$ as the input $T^{E}_{p}$ of the generator using the learnable mask token,
$T^{G}=\{t^{g}_{i}\}_{i\in\mathcal{G}}=\text{Generator}(T^{E}_{p},C^{s}_{RE},\mathcal{C}),~{}T^{G}\in\mathbb{R}^{G\times D},$ (11)
where $C^{s}_{RE}$ is the relative position encoding of the sorted center points $C^{s}$ to adapt to the needs of the generation task, and $\mathcal{C}$ is optional conditional information (e.g., category, image, and text).
Reconstructive Training.
Our training target is to reconstruct the masked tokens along with the center point coordinates $\hat{c}_{i}$. The reconstruction of center points serves the generation task because our VQVAE decoder requires absolute position encoding as input to reconstruct 3D shapes. Our loss function consists of two parts: the cross-entropy loss for token reconstruction and the mean square error loss for center point reconstruction, which can be expressed as:
$\mathcal{L}_{rt}=-\mathbb{E}_{t^{q}_{i}\in\mathcal{Q}}[\mathop{\sum}\limits_{\forall i,m_{i}=1}\log p(t^{q}_{i}|T^{G}_{m})]+\frac{1}{|M|}\mathop{\sum}\limits_{\forall i,m_{i}=1}\|c_{i},\hat{c}_{i}\|^{2}$ (12)
Where $T^{G}_{m}$ represents the masked portion of the tokens outputted by the generator.
3.4 Iterative Generation
After pretraining, following Li_2023_CVPR , we employ an iterative generation strategy to generate diverse 3D shapes. The generation process starts with all tokens being considered as masked tokens. Using a cosine function, we determine the number of masked tokens to be replaced in each iteration. Specifically, in each iteration, based on the predicted probabilities from the generator, we sample and replace those tokens with higher probabilities.
Table 1: Shape classification results on ModelNet40 Wu_2015_CVPR and ScanObjectNN Uy_2019_ICCV . We report the accuracy (%) of both training from scratch and fine-tuning models on ModelNet40 and ScanObjectNN datasets.
Methods Reference ModelNet40 ScanObjectNN
OBJ_BG OBJ_ONLY PB_T50_RS
training from scratch
PointNet Qi_2017_CVPR CVPR 2017 89.2 73.3 79.2 68.0
PointNet++ qi2017pointnet++ Neurips 2017 90.5 82.3 84.3 77.9
PointCNN li2018pointcnn Neurips 2018 92.2 86.1 85.5 78.5
DGCNN wang2019dynamic TOG 2019 92.9 82.8 86.2 78.1
GBNet 9410405 TMM 2021 93.8 - - -
MVTN Hamdi_2021_ICCV ICCV 2021 93.8 92.6 - 82.8
PointNeXt qian2022pointnext Neurips 2022 94.0 - - -
fine-tuning
Point-BERT Yu_2022_CVPR CVPR 2022 93.2
Mask-Point liu2022masked ECCV 2022 93.8 89.3 88.1 84.3
Point-MAE pang2022masked ECCV 2022 93.8 90.0 88.2 85.2
Point-M2AE zhang2022point Neurips 2022 94.0 91.2 88.8 86.4
PointGPT NEURIPS2023_5ed5c3c8 Neurips 2023 94.0 91.6 90.0 86.9
Point-MAGE - 94.2 92.9 91.0 87.1
Table 2: Few-shot classification on ModelNet40Wu_2015_CVPR . We report the average accuracy (%) and standard deviation (%) of 10 independent experiments.
Methods Reference 5-way 10-way
10-shot 20-shot 10-shot 20-shot
training from scratch
DGCNN wang2019dynamic TOG 2019 31.6 ± 2.8 40.8 ± 4.6 19.9 ± 2.1 16.9 ± 1.5
OcCo Wang_2021_ICCV ICCV 2021 90.6 ± 2.8 92.5 ± 1.9 82.9 ± 1.3 86.5 ± 2.2
fine-tuning
Transformer + OcCo Yu_2022_CVPR CVPR 2022 94.0 ± 3.6 95.9 ± 2.3 89.4 ± 5.1 92.4 ± 4.6
Point-BERT Yu_2022_CVPR CVPR 2022 94.6 ± 3.1 96.3 ± 2.7 91.0 ± 5.4 92.7 ± 5.1
Mask-Point liu2022masked ECCV 2022 95.0 ± 3.7 97.2 ± 1.7 91.4 ± 4.0 93.4 ± 3.5
Point-MAE pang2022masked ECCV 2022 96.3 ± 2.5 97.8 ± 1.8 92.6 ± 4.1 95.0 ± 3.0
Point-M2AE zhang2022point Neurips 2022 96.8 ± 1.8 98.3 ± 1.4 92.3 ± 4.5 95.0 ± 3.0
PointGPT NEURIPS2023_5ed5c3c8 Neurips 2023 96.8 ± 2.0 98.6 ± 1.1 92.6 ± 4.6 95.2 ± 3.4
Point-MAGE - 97.2 ± 1.9 98.6 ± 0.7 92.7 ± 4.0 95.4 ± 3.0
4 Experiments
In this section, we first provide a detailed overview of the pretraining setup of Point-MAGE. We then evaluate the pretrained model’s performance in both representation learning and generation capabilities. Finally, to gain a deeper understanding of the roles and importance of each component in the model, we conduct ablation studies on Point-MAGE. More results can be found in the Appendix.
4.1 Pre-training Setups
Point-MAGE is pre-trained on the ShapeNet chang2015shapenet dataset, which includes 55 different object categories and over 50,000 3D models. From the surface of each 3D shape, 2048 points are uniformly sampled and further divided into 512 patches, each containing 32 points. In the first stage, we pre-trained the VQVAE to obtains a discrete token for each point patch, resulting in a token sequence of 512 tokens. In the second stage, we determine the variable masking ratio with $\mu=0.55$ and $\sigma=0.25$ and pre-train the extractor-generator architecture based on masked token reconstruction. Additionally, to validate the capability of integrating the Point-MAGE framework with advanced model designs, we pre-trained a multi-scale extractor model following zhang2022point and evaluated its representation learning performance across various downstream tasks. We used the AdamW optimizer for pre-training the two stages for 800 and 300 epochs, respectively. Further implementation details and hyperparameter settings are provided in the Appendix.
4.2 Representation Learning
To demonstrate the representation learning performance of our method, we conducted experiments on shape classification, few-shot learning, and part segmentation. Additionally, the Appendix will provide further research and discussion on the properties of Point-MAGE.
Shape Classification.
We conducted shape classification fine-tuning on two datasets. The ModelNet40 Wu_2015_CVPR dataset includes 3D CAD models of 40 categories, and we followed the official data split provided in  Wu_2015_CVPR for our experiments. As shown in Table 1, Point-MAGE achieved a classification accuracy of 94.2% on this dataset, surpassing other existing pre-trained models. The ScanObjectNN  Uy_2019_ICCV dataset, derived from real-world indoor scans, contains approximately 15,000 objects across 15 categories. We evaluated this dataset using three different data splits: OBJ-BG, OBJ-ONLY, and PB-T50-RS. The experimental results indicate that Point-MAGE consistently outperformed other models across all splits. Notably, in the OBJ-BG setting, Point-MAGE achieved an accuracy of 92.9%, outperforming the second-best pre-trained model NEURIPS2023_5ed5c3c8 by 1.3 percentage points.
Few-shot Classification.
We evaluated the few-shot learning performance of Point-MAGE by conducting N-way, K-shot experiments on the ModelNet40Wu_2015_CVPR dataset. Specifically, we randomly selected N classes from ModelNet40 and sampled K objects from each class. Each experiment was repeated 10 times, and the average and standard deviation of the results were calculated. We evaluated each of the four combinations of settings with N = {5, 10} and K = {10, 20}. As shown in the table 2, we achieved SOTA performance in all of them, demonstrating that Point-MAGE’s pre-trained 3D representation is well-suited for few-shot learning tasks.
Table 3: Part segmentation on ShapeNetPart yi2016scalable . We report the mean IoU (%) of all part categories (’$\text{mIoU}_{C}$’) and the mean IoU of all instances (’$\text{mIoU}_{I}$’) in the dataset.
       Methods        Reference        $\text{mIoU}_{C}$uparrow        $\text{mIoU}_{I}$uparrow
       training from scratch
       PointNet Qi_2017_CVPR        CVPR 2017        80.4        83.7
       PointNet++ qi2017pointnet++        Neurips 2017        81.9        85.1
       DGCNN wang2019dynamic        TOG 2019        82.3        85.2
       fine-tuning
       Transformer + OcCo Yu_2022_CVPR        CVPR 2022        83.4        85.1
       Point-BERT Yu_2022_CVPR        CVPR 2022        84.1        85.6
       Mask-Point liu2022masked        ECCV 2022        84.4        86.0
       Point-MAE pang2022masked        ECCV 2022        -        86.1
       Point-M2AE zhang2022point        Neurips 2022        84.9        86.5
       PointGPT NEURIPS2023_5ed5c3c8        Neurips 2023        84.1        86.2
       Point-MAGE        -        85.0        86.5
Table 4: Unconditional generation. We report four metrics for evaluating the unconditional shape generation task.
Methods Surface-FPD $\downarrow$ Surface-KPD (×$10^{3}$$\downarrow$ Rendering-FID  $\downarrow$ Rendering-KID (×$10^{3}$$\downarrow$
Grid-$8^{3}$ zhang20223dilg 4.03 6.15 32.78 14.12
3DILG zhang20223dilg 1.89 2.17 24.83 10.51
Point-MAGE 0.98 1.34 22.79 11.48
Figure 2: Visualization of unconditional generation. All models were trained on the complete ShapeNet dataset, and we sampled nine generated results for rendering for each model.
Part Segmentation.
Table 3 presents the results of average IoU for all categories ($\text{mIoU}{C}$) and all instances ($\text{mIoU}{I}$) in the part segmentation experiments on the ShapeNetPart yi2016scalable dataset. Compared to the fully supervised method DGCNNwang2019dynamic , Point-MAGE achieved significant improvements of 2.7% and 1.3% in these two key evaluation metrics, respectively. Our method also outperforms other pretrained models, achieving a $\text{mIoU}_{C}$ of 85.0%, which ranks first.
4.3 3D Shape Generation
Unconditional Generation.
After completing the pre-training stage, Point-MAGE is capable of performing unconditional generation tasks. Fréchet Inception Distance (FID) and Kernel Inception Distance (KID) are commonly used to evaluate the generation quality of generative models. Following zhang20233dshape2vecset , we use Rendering-FID and Rendering-KID metrics to evaluate the FID and KID of rendered images of 3D shapes, as well as Surface-FPD and Surface-KPD metrics to evaluate the FID and KID of sampled point clouds from 3D shapes. Table 4 compares the performance of various methods in unconditional shape generation on ShapeNet. 3DILG also adopts a two-stage generation strategy and completes the generation training through an autoregressive process. Point-MAGE outperforms other methods on all three evaluation metrics, indicating that our method is capable of generating 3D shapes that match the real distribution. Additionally, we visualize the rendered results of generated shapes for all methods in Figure 2.
Figure 3: Visualization of category-conditional generation. We selected three category labels (airplane, faucet, and table) to display the rendered results of the generated shapes.
Figure 4: Visualization of masked point clouds conditional generation. We used a block masking strategy to mask 70% of the complete point cloud as conditional input. The generated results demonstrated the ability of different models to reconstruct the complete shapes.
Conditional Generation.
We evaluated the conditional generation capabilities of Point-MAGE under two different settings. First, we retrained the model using the 54 categories from the ShapeNet dataset as conditional embeddings. Figure 3 displays the rendered results of generated shapes for three categories (airplane, faucet, and table), with nine instances shown for each category. These visualizations demonstrate our method’s ability to generate realistic 3D shapes based on categorical conditions. Next, we further evaluated Point-MAGE against other SSL models in the masked generation task. As shown in Figure 4, these models typically train by reconstructing local point clouds, which primarily allows them to complete point cloud reconstruction tasks and depends on the provision of the central point positions. This approach, to some extent, leaks complete shape information to the decoder. In contrast, our method can reconstruct 3D shape meshes using the minimal amount of point cloud information and does not rely on masked center point information during the generation process.
4.4 Ablation Studies
In this section, we discuss three aspects that influence the representation learning and generation capabilities of the pre-trained model: the transformer design of the extractor-generator architecture, the impact of different inputs, and the effect of varying masking ratios.
Transformer Depth.
As shown in Table 5(a), we used the standard ViT architecture to build the extractor-generator structure and explored four different transformer depth combinations. The results indicate that increasing the model depth significantly enhances the performance of the pre-trained model. Additionally, consistent with trends observed in previous masked modeling studies he2022masked ; pang2022masked ; zhang2022point , an asymmetric extractor-generator structure is more beneficial for representation learning. This reliance on the extractor to obtain high-quality representations from a small number of visible tokens is due to the use of a relatively simple generator to complete the reconstruction task.
Input Type.
In Table 5(b), we investigate the impact of two different input types on model performance. Specifically, after vector quantizing the initial features of the point cloud patches, we consider using either the quantized features or their indices in the codebook as inputs. Experimental results show that using index-embedded tokens benefits the generation task, but its generalization ability across datasets is limited, which is detrimental to fine-tuning on downstream tasks, leading to poorer representation learning capabilities. This contrasts sharply with the results shown by MAGE Li_2023_CVPR , where using index embeddings performs well in both aspects. We believe this characteristic of MAGE arises from its pre-training on a larger-scale dataset, and the advantages of image data over point cloud data in terms of regularity and consistency, which reduce the impact of different data sources.
Masking Design.
To determine the optimal masking ratio, in Table 5(c), we compare the impact of various masking ratio distributions on fine-tuning and near-unconditional generation task results. Here, $\mu$ and $\sigma$ represent the mean and standard deviation of the truncated Gaussian distribution, respectively. When $\sigma=0$, a fixed masking ratio is used. In this case, when $\mu=1.0$, all tokens are masked, preventing the extractor from obtaining useful information, making the fine-tuning process essentially equivalent to training from scratch. In other cases, since the generator does not experience the process of reconstructing from zero, it fails to complete the generation task, and the generation results are marked as not applicable (N/A). The findings indicate that employing a variable masking ratio not only enables the completion of generation tasks but also encourages the representation learning process to capture superior features.
Table 5: Ablation experiments. The fine-tuned accuracy on the ScanObjectNN Uy_2019_ICCV (OBJ_BG) and the Rendering-FID (R-FID) for unconditional generation are reported.
(a) Transformer depth
Extractor Generator Acc (%) R-FID $\downarrow$
6 6 91.1 29.07
8 6 91.7 27.85
12 8 92.4 22.79
12 12 91.9 20.94
(b) Input type
Type Acc (%) R-FID $\downarrow$
Quantized features 92.4 22.79
Index-embedded tokens 87.3 22.36
(c) Masking ratio distribution
$\mu=1.0$ $\mu=0.7$ $\mu=0.6$ $\mu=0.55$ $\mu=0.5$ $\mu=0.55$ $\mu=0.55$ $\mu=0.55$
$\sigma=0$ $\sigma=0$ $\sigma=0$ $\sigma=0$ $\sigma=0$ $\sigma=0.15$ $\sigma=0.25$ $\sigma=0.5$
Acc (%) 86.2 91.8 91.9 92.0 91.8 92.2 92.4 92.1
R-FID $\downarrow$ 26.72 N/A N/A N/A N/A 23.58 22.79 23.65
5 Conclusion
We propose Point-MAGE, a novel unified framework for shape generation and representation learning on point clouds. Our approach serializes point cloud segments and employs a two-stage process to obtain quantized tokens, applying a variable masking ratio to accommodate both tasks simultaneously. We have achieved extensive results in shape classification, few-shot transfer learning, unconditional shape generation, and conditional generation. In terms of representation learning, our method has set a new SOTA result. In the future, we plan to extend Point-MAGE to pretrain on larger datasets to achieve even better performance. We hope our work draws the attention of researchers to the synergistic effects of combining two types of models in 3D vision.
Appendix
Appendix A Implementation Details
Our pre-trained model was trained on four A800 80G GPUs, and the fine-tuned model was trained on two 3090 24G GPUs. Below, we provide detailed descriptions of the training procedures and hyperparameter settings for each task.
Positional Encodings.
For absolute position encoding, we use a two-layer multilayer perceptron (MLP) to encode the center point of each point cloud patch into a D-dimensional vector. For relative position encoding, we use a learnable embedding layer to embed the indices of the fixed-length point patch sequence.
Pre-training.
In the ShapeNet [4] dataset, we randomly sampled each 3D shape to generate a point set of 2048 points, which served as input data for the two-stage model training. To improve the model’s generalization ability, we employed data augmentation techniques, including random scaling and flipping of the sampled point clouds. In the first stage, the point set was divided into 512 point patches, each containing 32 points. The VQVAE’s encoder and decoder have a ViT model depth of 6. The training parameters were set as follows: a batch size of 128, an initial learning rate of $5\times 10^{-4}$, and a weight decay coefficient of $5\times 10^{-2}$. In the second stage, the point patches were tokenized to obtain a token sequence with a length of 512. The extractor’s depth was set to 12, and the generator’s depth was set to 8. Additionally, we pre-trained a multi-scale extractor with a scale value of 3, where the number of point patches for each scale was [512, 256, 64], and the corresponding token dimensions were [96, 192, 384]. The training parameters for this stage were set as follows: a batch size of 256, an initial learning rate of $10^{-3}$, and a weight decay coefficient of $5\times 10^{-2}$. In both stages, we used a cosine annealing learning rate scheduler and employed a warm-up phase during the first 10 epochs of training.
Shape Classification.
For the shape classification task, we finetuned the Point-MAGE model on two datasets: ModelNet40[45] and ScanObjectNN[36]. The ModelNet40[45] dataset consists of synthetic 3D shapes from 40 categories. We used 9843 samples for training and 2468 samples for validation. For each sample, we sampled 2048 points to construct the input for the model. On the other hand, the ScanObjectNN[36] dataset contains real-world 3D shapes from 15 categories, divided into three subsets: OBJ-BG, OBJ-ONLY, and PB-T50-RS, with a total of 11416 training point clouds and 2882 validation point clouds. For these point cloud samples, we sampled 2048 points as input. During the finetuning training, both datasets were trained for 300 epochs, with a batch size of 32, learning rate of $5\times 10^{-4}$, and weight decay coefficient of $5\times 10^{-2}$. We kept the other training hyperparameters consistent with the settings from the pretraining stage.
Few-shot Classification.
We randomly selected K categories from the 40 categories in the ModelNet40[45] dataset and sampled N + 20 3D shapes for each selected category, where N shapes were used for model training and the remaining 20 shapes were used for testing. We designed four different few-shot learning settings: 5-way 10-shot, 5-way 20-shot, 10-way 10-shot, and 10-way 20-shot. Except for the training epochs set to 150, other training settings remained consistent with the shape classification experiments.
Part Segmentation.
We conducted experiments on the ShapeNetPart[53] dataset, which includes 16 object categories and 50 part categories, totaling 16,881 3D shape models. We selected 14,007 shapes from the dataset for training and reserved 2,874 shapes for validation. During training, the model was optimized for 300 epochs, with each batch processing 16 input point clouds, each containing 2048 points. The learning rate was set to $2\times 10^{-4}$, and the weight decay coefficient was 0.1.
Appendix B Additional Results and Analysis
Learning Curves.
As shown in Figure 5, we compared the learning curves of our method and Point-M2AE [57] during the pre-training phase for reconstructing each local point cloud patch. We used the same single-layer convolutional reconstruction head as Point-M2AE for reconstructing each point cloud patch. It is noteworthy that our pre-training process optimizes the token reconstruction loss rather than the traditional point cloud Chamfer distance loss. Surprisingly, even without using the corresponding loss function, our method can still effectively learn to reconstruct local point cloud patches by reconstructing semantic tokens. This finding reveals that the feature information learned through point cloud reconstruction is actually a subset of the feature information learned by our method. Ultimately, we showed the difference in loss values near convergence: our method’s reconstruction loss is slightly higher than that of Point-M2AE. We believe this difference is due to the point cloud sampling variance issue discussed in the main text. Using the Chamfer distance loss forces the model to learn specific surface sampling, which may lead to a lower loss value upon convergence but does not benefit the model in learning the general characteristics of surface shapes.
Figure 5: Visualization of masked point clouds conditional generation.
Performance of Linear SVM.
As shown in Table 6, we provide detailed reports on the linear SVM classification accuracy of Point-MAGE and other SSL methods on the ModelNet40[45] dataset, and find that our method achieves outstanding performance. Specifically, compared to the pioneering method Point-BERT[54] for point cloud masking SSL, our model achieves a significant improvement of 5.7% in classification accuracy.
Table 6: Linear SVM Performance on ModelNet40[45]. We report the shape classification accuracy (%).
Methods Reference Acc (%)
3D-GAN[44] Neurips 2016 83.3
Latent-GAN[38] TMM 2020 85.7
FoldingNet[52] CVPR 2018 88.4
MAP-VAE[14] ICCV 2019 88.4
VIP-GAN[13] AAAI 2019 90.2
DGCNN + Jiasaw[33] Neurips 2019 90.6
DGCNN + OcCo[41] ICCV 2021 90.7
DGCNN + CrossPoint[2] CVPR 2022 91.2
Transformer + OcCo[54] CVPR 2022 89.6
Point-BERT[54] CVPR 2022 87.4
Point-MAE[25] ECCV 2022 91.0
Point-M2AE[57] Neurips 2022 92.9
Point-MAGE - 93.1
t-SNE Visualizations.
In Figure 6, we applied t-SNE to visualize the feature distributions on the ModelNet40 [45] dataset, showing the results of three different training strategies: (a) training from scratch, (b) pre-training, and (c) fine-tuning. The observations reveal that when the Point-MAGE model is trained from scratch, there is significant overlap of features from different categories in the feature space. However, after pre-training, the model can generate more distinguishable features for different categories, enhancing feature recognizability. Furthermore, with fine-tuning, we found that the features of samples within the same category become more tightly clustered in the feature space, while the features of samples from different categories show clear separation. This indicates that the fine-tuning process further optimizes the clustering of features and enhances the model’s ability to distinguish between features of different categories.
Figure 6: t-SNE visualization on ModelNet40[45].
Part Segmentation.
In the part segmentation task on the ShapeNetPart dataset [53], the Point-MAGE model demonstrated outstanding performance. As shown in Table 7, Point-MAGE achieved the highest accuracy in 10 out of 16 categories. Overall, the average accuracy of Point-MAGE was the best among all methods, proving its effectiveness in handling complex 3D shapes and part segmentation tasks.
Table 7: Additional results of part segmentation on ShapeNetPart[53]. We report the IoU (%) for each category.
Methods aero bag cap car chair earp. guit. knif. lamp lapt. mot. mug pist. rock. skt. table
PointNet[26] 83.4 78.7 82.5 74.9 89.6 73.0 91.5 85.9 80.8 95.3 65.2 93.0 81.2 57.9 72.8 80.6
PointNet++[27] 82.4 79.0 87.7 77.3 90.8 71.8 91.0 85.9 83.7 95.3 71.6 94.1 81.3 58.7 76.4 82.6
DGCNN[42] 84.0 83.4 86.7 77.8 90.6 74.7 91.2 87.5 82.8 95.7 66.3 94.9 81.1 63.5 74.5 82.6
Point-BERT[54] 84.3 84.8 88.0 79.8 91.0 81.7 91.6 87.9 85.2 95.6 75.6 94.7 84.3 63.4 76.3 81.5
Mask-Point[21] 84.2 85.6 88.1 80.3 91.2 79.5 91.9 87.8 86.2 95.3 76.9 95.0 85.3 64.4 76.9 81.8
Point-MAE[25] 84.3 85.0 88.3 80.5 91.3 78.5 92.1 87.4 86.1 96.1 75.2 94.6 84.7 63.5 77.1 82.4
Point-MAGE 85.2 87.2 89.2 81.4 91.8 80.4 92.4 88.4 86.0 96.0 78.0 95.3 85.0 64.1 77.6 82.0
Appendix C Limitations and Future Directions
Despite the significant enhancements our method brought to the capabilities of the pre-trained model, these improvements did not translate into substantial performance gains across all downstream tasks. For instance, in the part segmentation experiments, the observed performance gain was relatively small. Additionally, the performance of our method was influenced by the input type (quantized features or index embeddings), and it did not fully replicate the success of MAGE [19] in achieving optimal representation learning and generative performance simultaneously in the 2D image domain. We attribute this phenomenon partly to the fact that the scale of the point cloud pre-training dataset is much smaller than that of the 2D image domain. Therefore, we expect that by expanding the pre-training dataset, we can further improve model performance and achieve a better unification of representation learning and generative performance.

### v2

@@
 Abstract
-In the field of 2D image generation modeling and representation learning, Masked Generative Encoder (MAGE) has demonstrated the synergistic potential between generative modeling and representation learning.
-Inspired by this, we propose Point-MAGE to extend this concept to point cloud data.
-Specifically, this framework first utilizes a Vector Quantized Variational Autoencoder (VQVAE) to reconstruct a neural field representation of 3D shapes, thereby learning discrete semantic features of point patches.
-Subsequently, by combining the masking model with variable masking ratios, we achieve synchronous training for both generation and representation learning.
-Furthermore, our framework seamlessly integrates with existing point cloud self-supervised learning (SSL) models, thereby enhancing their performance.
-We extensively evaluate the representation learning and generation capabilities of Point-MAGE.
-In shape classification tasks, Point-MAGE achieved an accuracy of 94.2% on the ModelNet40 dataset and 92.9% (+1.3%) on the ScanObjectNN dataset.
-Additionally, it achieved new state-of-the-art performance in few-shot learning and part segmentation tasks.
-Experimental results also confirmed that Point-MAGE can generate detailed and high-quality 3D shapes in both unconditional and conditional settings.
+Representation and generative learning, as reconstruction-based methods, have demonstrated their potential for mutual reinforcement across various domains.
+In the field of point cloud processing, although existing studies have adopted training strategies from generative models to enhance representational capabilities, these methods are limited by their inability to genuinely generate 3D shapes.
+To explore the benefits of deeply integrating 3D representation learning and generative learning, we propose an innovative framework called Point-MGE.
+Specifically, this framework first utilizes a vector quantized variational autoencoder to reconstruct a neural field representation of 3D shapes, thereby learning discrete semantic features of point patches.
+Subsequently, we design a sliding masking ratios to smooth the transition from representation learning to generative learning.
+Moreover, our method demonstrates strong generalization capability in learning high-capacity models, achieving new state-of-the-art performance across multiple downstream tasks.
+In shape classification, Point-MGE achieved an accuracy of 94.2% (+1.0%) on the ModelNet40 dataset and 92.9% (+5.5%) on the ScanObjectNN dataset.
+Experimental results also confirmed that Point-MGE can generate high-quality 3D shapes in both unconditional and conditional settings.
 1 Introduction
-Recent studies have confirmed that representation learning frameworks based on masked modeling Yu_2022_CVPR ; liu2022masked ; pang2022masked ; zhang2022point ; NEURIPS2023_5ed5c3c8 can extract high-quality 3D representations.
-At the same time, generative models zhang20223dilg ; Mittal_2022_CVPR ; Yan_2022_CVPR ; zhang20233dshape2vecset have also made significant achievements in generating realistic 3D shapes.
-However, there is little research that attempts to combine the potential of these two types of models to explore their joint application in 3D shapes.
-We note that in the field of 2D images, Masked Generative Encoder (MAGE) Li_2023_CVPR proposed an innovative unified framework for generation training and representation learning, and experimentally verified the synergistic effects between these two tasks.
-This inspired us to pose a question: Can the characteristics demonstrated in the MAGE framework be extended to point cloud data to achieve deeper 3D representation learning and generation capabilities?
-The strategy of MAGE Li_2023_CVPR is to combine Masked Image Modeling (MIM) with variable masking ratios.
-Specifically, this method conducts generation training under high masking ratios, while representation learning is performed under low masking ratios.
-Following this idea, we explore combining masked point cloud modeling with variable masking ratios to realize the Point-MAGE framework.
-However, due to the fundamental differences in structure and distribution between images and point cloud data, this direct transfer faces several challenges.
-Firstly, MAGE Li_2023_CVPR effectively obtains discrete semantic features from images through pixel-wise reconstruction training.
-However, directly using point clouds as reconstruction targets introduces sampling bias.
-This is because 3D shapes are inherently continuous, while point clouds obtained through sampling merely represent them discretely.
-This leads to non-zero loss between different point cloud samples of the same object, affecting the accurate understanding of shapes.
-Moreover, methods solely relying on point cloud reconstruction fail to acquire more advanced representations of 3D shapes, such as neural radiance fields or triangle meshes, thereby limiting the potential of models in generation tasks.
-Therefore, simply applying variable masking ratios to existing point cloud SSL methods Yu_2022_CVPR ; pang2022masked ; zhang2022point ; NEURIPS2023_5ed5c3c8 cannot achieve the goals of Point-MAGE, as these existing methods typically use Chamfer distance (CD) as the reconstruction loss for point cloud reconstruction.
-To address these challenges, we instead focus the reconstruction targets on 3D shapes represented as neural radiance fields (NeRF) and extract discrete semantic features of point clouds through a VQVAE van2017neural .
-Point-MAGE leverages these high-level semantic discrete features as input and reconstruction targets, learning the probability distribution of masked tokens.
-After training, our method not only iteratively performs three-dimensional shape generation tasks but also resolves sampling bias in representation learning through the vector quantization process.
-This is because different point cloud samples from the same local block are unified into the same discrete token during the vector quantization stage, ensuring consistency of representations.
+Recent studies have confirmed that masked modeling-based representation learning frameworks Yu et al.
+(2022); Pang et al.
+(2022); Zhang et al.
+(2022b) can extract high-quality 3D representations.
+At the same time, generative modelss Zhang et al.
+(2022a); Mittal et al.
+(2022); Zhang et al.
+(2023a) have achieved remarkable success in generating realistic 3D shapes.
+Nevertheless, few studies have explored the deep integration of these two models to unlock their combined potential in 3D shape processing.
+In natural language processing (NLP) Touvron et al.
+(2023) and 2D image domains Li et al.
+(2023), we have witnessed the versatility of pre-trained models: not only can they be used for diverse generative tasks such as question answering (QA), code generation, and image generation, but they can also serve as feature extractors for various downstream tasks through fine-tuning.
+These achievements prompt us to consider: can the benefits of this model integration be similarly applied to the representation of point cloud data and 3D shape generation?
+To this end, we conducted an in-depth analysis of the characteristics of both models and found that masked reconstruction methods offer a potential pathway to achieve this integration, though several challenges remain.
+Firstly, generative learning is easily trained on discrete distributions, whereas existing point cloud representation learning Pang et al.
+(2022); Zhang et al.
+(2022b) tends to operate within continuous feature spaces.
+In prior work, Point-Bert Yu et al.
+(2022) learns a tokenizer by reconstructing the input point cloud and optimizing the chamfer distance (CD) loss.
+However, using point clouds directly as the reconstruction target introduces a sampling bias.
+This is because 3D shapes are inherently continuous, but point clouds sampled from these shapes can only approximate them in a discrete form.
+This sampling bias can lead to non-zero loss differences between different point cloud samples of the same object, thereby affecting the precise understanding of shape features.
+Moreover, methods that rely solely on point cloud reconstruction are limited in capturing high-level representations of 3D shapes.
+For instance, they cannot generate richer shape descriptions such as neural radiance fields (NeRFs) or triangular meshes, which restricts the model’s potential in generative tasks.
+To address this, we shift the reconstruction target to 3D shapes represented by NeRFs and utilize a vector quantized variational autoencoder (VQVAE) Van Den Oord et al.
+(2017) to extract discrete semantic features from point clouds.
+The Point-MGE framework leverages these high-level semantic discrete features not only as input but also as reconstruction targets, thereby learning the probability distribution of masked tokens.
+During the vector quantization process, different point cloud samples from the same local region are consistently mapped to the same discrete tokens, ensuring consistency and robustness in the representations.
 Secondly, the inherent unordered nature of point cloud poses a challenge for Transformer-based generative models, as these models require serialized inputs.
 To address this issue, we partition the complete point cloud into multiple patches, each defined by a fixed number of points and a central point.
-To serialize these patches, we adopt the Morton sorting algorithm used in PointGPT NEURIPS2023_5ed5c3c8 , which can sort point cloud patches while preserving geometric relationships.
-In generation task, since the central point of point cloud patch is unknown, we introduce relative position encoding in the generator of Point-MAGE as a replacement for traditional absolute position encoding, and use the central points as the reconstruction target.
-Finally, we conducted an analysis of the synergistic effects of the two tasks.
-We found that generative models benefit from the variable masking ratio.
-Unlike traditional methods that reconstruct all tokens from scratch, our approach leverages a small number of visible tokens to provide additional sample information, which not only accelerates convergence but also enhances the model’s final performance.
-For representation learning, by shifting the reconstruction target from raw point clouds to discrete semantic features, our framework successfully overcomes the variance issue in point cloud sampling and focuses more on the global features of 3D shapes.
-Furthermore, the Point-MAGE framework can be integrated with SSL model designs proposed in previous work, further enhancing the performance of representation learning.
-Experimental results demonstrate the effectiveness of this strategy, with Point-MAGE achieving state-of-the-art (SOTA) performance on multiple downstream tasks.
-Specifically, our method achieved an accuracy of 94.2% on the ModelNet40 dataset Wu_2015_CVPR and 92.9% on the ScanObjectNN dataset yi2016scalable , respectively, for shape classification tasks.
-In the part segmentation task on the ShapeNetPart yi2016scalable dataset, our method achieved a mean class Intersection over Union (IoU) of 85.0%.
-Additionally, we demonstrated the pre-trained model’s ability to generate realistic 3D shapes in both unconditional and conditional settings, further showcasing the flexibility and powerful generative capabilities of Point-MAGE.
-Our main contributions can be summarized as follows: (I) We propose Point-MAGE, a unified framework for point cloud data representation learning and generative training that combines a two-stage training method with a variable masking ratio design.
-(II) Point-MAGE overcomes the variance issue in point cloud sampling and can be integrated with model architectures from previous SSL works, setting a new SOTA benchmark for downstream tasks such as shape classification, few-shot learning, and part segmentation.
-(III) Point-MAGE is capable of generating high-quality 3D shapes in both unconditional and conditional settings.
+To serialize these patches, we adopt the Morton sorting algorithm used in PointGPT Chen et al.
+(2023), which can sort point cloud patches while preserving geometric relationships.
+In generation task, since the central point of point cloud patch is unknown, we introduce relative position encoding in the generator of Point-MGE as a replacement for traditional absolute position encoding, and use the central points as the reconstruction target.
+Finally, generative training tends to adopt higher masking ratios, or even full masking, whereas representation learning typically uses relatively low masking ratios.
+In 2D image processing, MAGE Li et al.
+(2023) achieves variable masking ratios by randomly sampling them from a Gaussian distribution, utilizing these variable ratios to perform the task.
+However, due to the irregularity of point cloud data and the lack of support from large-scale datasets, such drastic fluctuations in masking ratios may adversely affect model convergence.
+To address this issue, we have carefully designed a sliding masking ratio mechanism to smoothly adjust the masking ratio during training.
+Specifically, the masking ratio gradually increases as training iterations progress, facilitating a smooth transition from representation learning to generative training.
+We also conducted a comparative analysis of the effects of different sliding curves.
+Point-MGE can scale to high-capacity model training and has demonstrated outstanding performance across multiple benchmark datasets.
+It achieved an accuracy of 94.2% on the ModelNet40 dataset Wu et al.
+(2015) for shape classification and 92.9% on the ScanObjectNN dataset Uy et al.
+(2019).
+Additionally, in the part segmentation task on the ShapeNetPart dataset Yi et al.
+(2016), our method attained an average class Intersection over Union (IoU) of 85.0%.
+Furthermore, we also showcased the capability of the pre-trained model to generate realistic 3D shapes in both unconditional and conditional settings.
+Our main contributions can be summarized as follows: (I) We propose Point-MGE, a unified framework for point cloud data representation learning and generative training that combines a two-stage training method with a sliding masking ratio design.
+(II) Point-MGE overcomes the variance issue in point cloud sampling and scales to high-capacity model training, setting a new SOTA benchmark for downstream tasks.
+(III) Point-MGE is capable of generating high-quality 3D shapes in both unconditional and conditional settings.
 2 Related Works
@@
 In recent years, self-supervised frameworks for point cloud representation learning have made significant progress, forming two main model paradigms.
-Firstly, methods based on contrastive learning Yang_2018_CVPR ; xie2020pointcontrast ; Navaneet_2020_CVPR ; Zhang_2021_ICCV ; Afham_2022_CVPR have played an important role in point cloud self-supervised learning (SSL).
+Firstly, methods based on contrastive learning Xie et al.
+(2020b); Zhang et al.
+(2021); Afham et al.
+(2022) have played an important role in point cloud self-supervised learning.
 These methods can explore the intrinsic properties of point clouds, but they often rely on cross-modal information or teacher models, which undoubtedly add extra burden to the training process.
-Secondly, more closely related to our work is another class of methods based on masked point cloud modeling.
-Inspired by the successful experience of masked modeling in natural language processing (NLP) radford2018improving ; devlin2018bert ; radford2019language ; brown2020language and image processing he2022masked ; Wei_2022_CVPR ; zhou2021ibot fields, recent research has focused on randomly masking point cloud regions and reconstructing these regions Yu_2022_CVPR ; pang2022masked ; zhang2022point ; NEURIPS2023_5ed5c3c8 , prompting the encoder to learn high-quality point cloud representations.
-In particular, Point-M2AE zhang2022point proposed an innovative hierarchical transformer structure, extending Point-MAE pang2022masked to capture multi-scale features of point clouds.
-On the other hand, PointGPT attempts to solve the problem of global shape leakage through autoregressive generation, but it does not have the ability to generate high-quality 3D models from scratch.
+Secondly, inspired by natural language processing Devlin et al.
+(2018); Brown et al.
+(2020) and 2D computer vision He et al.
+(2022); Wei et al.
+(2022), a more closely related area to our work is another category of methods based on masked point cloud modeling Yu et al.
+(2022); Pang et al.
+(2022); Zhang et al.
+(2022b); Chen et al.
+(2023).
+Point-M2AE Zhang et al.
+(2022b) proposed an innovative hierarchical transformer structure, extending Point-MAE Pang et al.
+(2022) to capture multi-scale features of point clouds.
+On the other hand, PointGPT Chen et al.
+(2023) attempts to solve the problem of global shape leakage through autoregressive generation, but it does not have the ability to generate high-quality 3D models from scratch.
 Additionally, methods based on masked region reconstruction are susceptible to point cloud sampling variance, which may introduce additional understanding and bias issues.
@@
 Generative Models for 3D Shapes.
-In the field of 3D generation, Generative Adversarial Networks (GANs) wu2016learning ; achlioptas2018learning ; Chen_2019_CVPR ; Ibing_2021_CVPR ; zheng2022sdf constitute a major generative paradigm, despite their limitations in training stability and mode collapse issues.
-Meanwhile, some research has turned to alternative approaches such as Normalizing Flows (NFs) dinh2014nice ; rezende2015variational ; yang2019pointflow , Energy-Based Models (EBMs) lecun2006tutorial ; xie2016theory ; Xie_2021_CVPR ; xie2020generative , and Autoregressive Models (ARs) Mittal_2022_CVPR ; cheng2022autoregressive ; Yan_2022_CVPR ; zhang20223dilg .
-Recently, inspired by the significant progress of Diffusion Models (DMs) dhariwal2021diffusion ; song2020score in image synthesis, some studies have begun to explore the application of diffusion models to 3D shape generation vahdat2022lion ; hui2022neural ; Shue_2023_CVPR ; Chou_2023_ICCV ; zhang20233dshape2vecset .
+In the field of 3D generation, generative adversarial networks (GANs) Achlioptas et al.
+(2018); Chen and Zhang (2019); Ibing et al.
+(2021); Zheng et al.
+(2022) constitute a major generative paradigm, despite their limitations in training stability and mode collapse issues.
+Meanwhile, some research has turned to alternative approaches such as normalizing flows (NFs) Dinh et al.
+(2014); Rezende and Mohamed (2015); Yang et al.
+(2019), energy-based models (EBMs) Xie et al.
+(2016, 2020a, 2021), and autoregressive models (ARs) Mittal et al.
+(2022); Cheng et al.
+(2022); Yan et al.
+(2022); Zhang et al.
+(2022a).
+Recently, inspired by the significant progress of diffusion models (DMs) Dhariwal and Nichol (2021) in image synthesis, some studies have begun to explore the application of diffusion models to 3D shape generation Zhou et al.
+(2021); Vahdat et al.
+(2022); Hui et al.
+(2022); Shue et al.
+(2023); Chou et al.
+(2023); Zhang et al.
+(2023a) .
 Although these models can generate realistic 3D models, they often lack the ability to extract high-quality semantic representations from 3D shapes, limiting their direct application in point cloud representation learning tasks.
 To address this challenge, we propose an innovative approach aimed at generating diverse samples while extracting high-quality semantic embeddings from 3D shapes to promote data generalization and enhance the performance of downstream tasks.
-We draw on the successful experience of MAGE Li_2023_CVPR in image synthesis and representation learning, and realize the organic combination of generation and representation learning on point clouds through a two-stage scheme.
-In the first stage, we use VQVAE van2017neural to train an encoder and decoder to obtain latent codes.
-Subsequently, in the second stage, we further train the extractor for representation learning through masked modeling and use a bidirectional transformer generator to complete the generation task.
 3 Method
-Point-MAGE is designed as a unified framework for both point cloud generation tasks and representation learning.
-As shown in Figure 1, the framework first divides the input point cloud into multiple local point blocks and serializes them based on their geometric positional relationships.
-Then, it uses a mini-PointNet Qi_2017_CVPR model for initial feature extraction of each point block.
-Subsequently, by training a VQVAE van2017neural model, these point block features are quantized into discrete semantic latents.
-Finally, the framework combines a variable masking ratio with a Vision Transformer (ViT)-based extractor-generator architecture to predict masked semantic tokens.
-Figure 1: Overall pipeline of Point-MAGE.
-The training process consists of two stages.
-In the first stage, a VQVAE van2017neural is used to reconstruct 3D shapes represented by NeRF, converting the input point cloud into a series of discrete semantic latents.
-The second stage is a masked token reconstruction process, where a ViT-based extractor-generator architecture is employed to extract high-quality feature representations from the unmasked tokens and to reconstruct the masked tokens.
+Point-MGE is designed as a unified framework for both 3D generation tasks and representation learning.
+As shown in Figure 1, the framework first divides the input point cloud into multiple local point patches and serializes them based on their geometric positional relationships.
+Then, it uses a mini-PointNet Qi et al.
+(2017a) model for initial feature extraction of each point patch.
+Subsequently, by training a VQVAE model, these features are quantized into discrete tokens.
+Finally, the framework combines a variable masking ratio with a vision transformer (ViT)-based extractor-generator architecture to predict masked tokens.
+Figure 1: Overall pipeline of Point-MGE.
+First, a VQVAE is used to reconstruct 3D shapes represented by NeRFs, converting the input point cloud into a series of discrete semantic tokens.
+Then, where a ViT-based extractor-generator architecture is employed to extract high-quality feature representations from the unmasked tokens and to reconstruct the masked tokens.
 3.1 Point Patch Embedding
@@
 Embedding.
-To provide serialized input, we follow NEURIPS2023_5ed5c3c8 and encode the coordinates of the central points $x_{i}$ into one-dimensional space using Morton ordering morton1966computer , and then sort them to obtain serialized central points $C^{s}\in\mathbb{R}^{G\times 3}$ and point cloud patches $P^{s}\in\mathbb{R}^{G\times 3}$.
-We project each patch into an embedding vector using the mini-PointNetQi_2017_CVPR module,
-[Uncaptioned image]${\color[rgb]{0.875,0.4765625,0.3671875}~{}:P^{s}_{E}}=\{e_{i}\}_{i\in\mathcal{G}}=\text{PointNet}\left(P^{s}\right),~{}P^{s}_{E}\in\mathbb{R}^{G\times D},$ (3)
+To provide serialized input, we follow Chen et al.
+(2023) and encode the coordinates of the central points $x_{i}$ into one-dimensional space using morton ordering Morton (1966), and then sort them to obtain serialized central points $C^{s}\in\mathbb{R}^{G\times 3}$ and point cloud patches $P^{s}\in\mathbb{R}^{G\times 3}$.
+We project each patch into an embedding vector using the mini-PointNetQi et al.
+(2017a) module,
+$P^{s}_{E}=\{e_{i}\}_{i\in\mathcal{G}}=\text{PointNet}\left(P^{s}\right),~{}P^{s}_{E}\in\mathbb{R}^{G\times D},$ (3)
 where $|\mathcal{G}|=G$, D is the embedding dimension of the point patch.
 3.2 Tokenization
-We train a VQVAE van2017neural to obtain a quantizer that embeds input point patches into a series of semantic latents.
+We train a VQVAE Van Den Oord et al.
+(2017) to obtain a quantizer that embeds input point patches into a series of semantic tokens.
 This tokenization step enables our model to focus on high-level semantics, which helps to overcome the variance issue in point cloud sampling and allows for a more comprehensive understanding of 3D shapes.
@@
 We use a ViT-based encoder to learn the local latents $T$.
-The input to the transformer includes the sorted point block embeddings $P^{s}_{E}$ and the sinusoidal positional encodings vaswani2017attention $C^{s}_{E}$ of the center point,
-$T=\{t_{i}\}_{i\in\mathcal{G}}=\text{Encoder}\left(P^{s}_{E},C^{s}_{E}\right),~{}T\in\mathbb{R}^{G\times D}.$ (4)
+The input to the transformer includes the sorted point block embeddings $P^{s}_{E}$ and the sinusoidal positional encodings Vaswani et al.
+(2017) $C^{s}_{E}$ of the center point,
+$T=\{t_{i}\}_{i\in\mathcal{G}}=\text{Enc}\left(P^{s}_{E},C^{s}_{E}\right),~{}T\in\mathbb{R}^{G\times D}.$ (4)
 The output of the encoder is then replaced with the closest vector in the codebook $\mathcal{Q}$ through vector quantization,
 $T^{Q}=\{t^{q}_{i}\}_{i\in\mathcal{G}}=\mathop{\arg\min}\limits_{t^{q}_{j}\in\mathcal{Q}}\|t^{q}_{j},t_{i}\|,~{}T^{Q}\in\mathbb{R}^{G\times D}.$ (5)
-NeRF Reconstruction.
-We utilize a ViT-based decoder along with a reconstruction head to reconstruct 3D shapes represented in NeRF.
-The decoder takes as input the quantized latents and the position encoding of the center points,
-$T^{D}=\{t^{d}_{i}\}_{i\in\mathcal{G}}=\text{Decoder}\left(T^{Q},C^{s}_{E}\right),~{}T^{D}\in\mathbb{R}^{G\times D}.$ (6)
+Reconstruction.
+We utilize a ViT-based decoder to reconstruct the 3D shape’s NeRFs $\hat{\mathcal{O}}(x)\in\{0,1\}$, where $\hat{\mathcal{O}}(x)=1$ indicates that the point $x$ in space belongs to the object.
+The decoder takes as input the quantized tokens and the position encoding of the center points,
+$T^{D}=\{t^{d}_{i}\}_{i\in\mathcal{G}}=\text{Dec}\left(T^{Q},C^{s}_{E}\right),~{}T^{D}\in\mathbb{R}^{G\times D}.$ (6)
 For any $x\in\mathbb{R}^{3}$, we use the nadaraya-watson estimator for interpolation to obtain its latent $t^{d}_{x}$.
 Then, we estimate $\mathcal{O}(x)$ through a multi-layer perceptron and a sigmoid activation function,
-$\mathcal{O}(x)=\text{Sigmoid}\left(\text{MLP}\left(x,t^{d}_{x}\right)\right),~{}\mathcal{O}(x)\in\left[0,1\right].$ (7)
+$\mathcal{O}(x)=\text{Sigmoid}\left(\text{MLP}\left(x,t^{d}_{x}\right)\right).$ (7)
 The optimization objective consists of binary cross-entropy (BCE) loss and an additional commitment loss,
-$\mathcal{L}_{n}=\mathbb{E}_{x\in\mathbb{R}^{3}}[\text{BCE}(\mathcal{O}(x),\hat{\mathcal{O}}(x))]+\lambda\mathbb{E}_{x\in\mathbb{R}^{3}}[\mathbb{E}_{i\in\mathcal{M}}\|\epsilon(t^{q}_{i})-t_{i}\|^{2}],$ (8)
-where $\hat{\mathcal{O}}(x)$ denotes the ground truth, and $\epsilon(\cdot)$ represents the stop-gradient operation. In this paper, $\lambda=0.25$ is a weighting coefficient.
+$\footnotesize\begin{split}\mathcal{L}_{n}=\mathbb{E}_{x\in\mathbb{R}^{3}}[\text{BCE}(\mathcal{O}(x),\hat{\mathcal{O}}(x))]+~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}\\ \lambda\mathbb{E}_{x\in\mathbb{R}^{3}}[\mathbb{E}_{i\in\mathcal{M}}\|\epsilon(t^{q}_{i})-t_{i}\|^{2}],\end{split}$ (8)
+where $\epsilon(\cdot)$ represents the stop-gradient operation. In this paper, $\lambda=0.25$ is a weighting coefficient.
 3.3 Pre-training
 After obtaining the tokenizer, we simultaneously train an extractor and a generator model through masking modeling.
-Below, we introduce the masking strategy, model design, and optimization objectives of Point-MAGE.
+Below, we introduce the masking strategy, model design, and optimization objectives of Point-MGE.
+Methods Reference Params(M) ModelNet40 ScanObjectNN
+w/o Vote w/ Vote OBJ_BG OBJ_ONLY PB_T50_RS
+Supervised Learning Only
+PointNet Qi et al. (2017a) CVPR 2017 3.5 - 89.2 73.3 79.2 68.0
+PointNet++ Qi et al. (2017b) Neurips 2017 1.5 - 90.7 82.3 84.3 77.9
+DGCNN Wang et al. (2019) TOG 2019 1.8 - 92.9 82.8 86.2 78.1
+PointMLP Ma et al. (2022) ICLR2022 12.6 94.1 94.5 - - 85.2
+P2P-HorNet Wang et al. (2022) NeurIPS 2022 195.8 94.0 - - - 89.3
+with Single-Modal Pre-Training
+Point-BERT Yu et al. (2022) CVPR 2022 22.1 92.7 93.2 87.4 88.1 83.1
+Mask-Point Liu et al. (2022) ECCV 2022 22.1 - 93.8 89.3 88.1 84.3
+Point-MAE Pang et al. (2022) ECCV 2022 22.1 93.2 93.8 90.0 88.2 85.2
+Point-M2AE Zhang et al. (2022b) Neurips 2022 15.3 93.4 94.0 91.2 88.8 86.4
+PointGPT Chen et al. (2023) Neurips 2023 19.5 - 94.0 91.6 90.0 86.9
+GPM Li et al. (2024) CVPR2024 41.5 94.0 94.3 90.2 90.0 84.8
+Point-MGE - 15.3 93.8 94.2 92.9 91.0 87.1
+Improvement (baseline: Point-BERT) - - +1.1 +1.0 +5.5 +2.9 +4.0
+Point-MAE-B Pang et al. (2022) ECCV 2022 120.1 - 94.2 94.2 93.9 90.2
+Point-M2AE-B Zhang et al. (2022b) Neurips 2022 77.5 - 94.3 95.2 94.3 91.2
+PointGPT-B Chen et al. (2023) Neurips 2023 82.1 - 94.4 95.8 95.2 91.9
+Point-MGE-B - 77.5 - 94.5 96.1 95.7 92.3
+with Cross-Modal Pre-Training
+ACT Dong et al. (2023) ICLR 2023 22.1 93.2 93.7 93.3 91.9 88.2
+Joint-MAE Guo et al. (2023) IJCAI 2023 - - 94.0 90.9 88.9 86.1
+I2P-MAE Zhang et al. (2023b) CVPR 2023 15.3 93.7 94.1 94.2 91.6 90.1
+Recon Qi et al. (2023) ICML 2023 44.3 94.1 94.5 95.2 93.3 90.6
+Table 1: Shape classification results on ModelNet40 Wu et al.
+(2015) and ScanObjectNN Uy et al.
+(2019).
+We report the accuracy (%) of both training from scratch and fine-tuning models on ModelNet40 and ScanObjectNN datasets.
 Masking Strategy.
 For a given point cloud $X\in\mathbb{R}^{n\times 3}$, we first use the trained tokenizer to obtain the corresponding discrete tokens $T^{Q}$.
-Subsequently, we introduce a variable mask ratio to apply masking, which can be calculated as:
-$m_{r}\sim N(\mu=0.55,\sigma^{2})~{}\text{with}~{}m_{r}\in\left[0.5,1\right]$ (9)
-Where $N$ represents the truncated Gaussian distribution.
-After masking, to reduce training time and memory consumption, we randomly drop a portion of masked tokens, ensuring that only $0.5\times G$ tokens are input to the extractor, following the MAGE Li_2023_CVPR .
+Subsequently, we introduce a sliding masking ratio to apply masking, which can be calculated as:
+$m_{r}=\beta+(1-\beta)(\frac{\gamma}{\Gamma})^{u},~{}m_{r}\in\left[\beta,1\right],$ (9)
+where, $\gamma$ and $\Gamma$ denote the current training iteration and the total number, respectively.
+The parameter $u>0$ is a hyperparameter used to control the trend of the sliding function.
+After masking, to reduce training time and memory consumption, we randomly drop a portion of masked tokens, ensuring that only $V=(1-\beta)\times G$ tokens are input to the extractor.
 Extractor and Generator.
@@
 The extractor processes the input tokens $T^{Q}_{m}$ and encodes them into the latent feature space,
-$T^{E}=\{t^{e}_{i}\}_{i\in\mathcal{G},m_{i}=0}=\text{Extractor}(T^{Q}_{m},C^{s}_{E}),~{}T^{E}\in\mathbb{R}^{\frac{G}{2}\times D},$ (10)
+$\footnotesize T^{E}=\{t^{e}_{i}\}_{i\in\mathcal{G},m_{i}=0}=\text{Ext}(T^{Q}_{m},C^{s}_{E}),~{}T^{E}\in\mathbb{R}^{V\times D},$ (10)
 where $M=[m_{i}]^{G}_{i=1}$ represents a binary mask indicating whether a token is masked or not.
 Then, the output of the extractor is padded to the complete length $G$ as the input $T^{E}_{p}$ of the generator using the learnable mask token,
-$T^{G}=\{t^{g}_{i}\}_{i\in\mathcal{G}}=\text{Generator}(T^{E}_{p},C^{s}_{RE},\mathcal{C}),~{}T^{G}\in\mathbb{R}^{G\times D},$ (11)
+$T^{G}=\{t^{g}_{i}\}_{i\in\mathcal{G}}=\text{Gen}(T^{E}_{p},C^{s}_{RE},\mathcal{C}),~{}T^{G}\in\mathbb{R}^{G\times D},$ (11)
 where $C^{s}_{RE}$ is the relative position encoding of the sorted center points $C^{s}$ to adapt to the needs of the generation task, and $\mathcal{C}$ is optional conditional information (e.g., category, image, and text).
@@
 Our loss function consists of two parts: the cross-entropy loss for token reconstruction and the mean square error loss for center point reconstruction, which can be expressed as:
-$\mathcal{L}_{rt}=-\mathbb{E}_{t^{q}_{i}\in\mathcal{Q}}[\mathop{\sum}\limits_{\forall i,m_{i}=1}\log p(t^{q}_{i}|T^{G}_{m})]+\frac{1}{|M|}\mathop{\sum}\limits_{\forall i,m_{i}=1}\|c_{i},\hat{c}_{i}\|^{2}$ (12)
+$\footnotesize\mathcal{L}_{rt}=-\mathbb{E}_{t^{q}_{i}\in\mathcal{Q}}[\mathop{\sum}\limits_{\forall i,m_{i}=1}\log p(t^{q}_{i}|T^{G}_{m})]+\frac{1}{|M|}\mathop{\sum}\limits_{\forall i,m_{i}=1}\|c_{i},\hat{c}_{i}\|^{2},$ (12)
 where $T^{G}_{m}$ represents the masked portion of the tokens outputted by the generator.
 3.4 Iterative Generation
-After pretraining, following Li_2023_CVPR , we employ an iterative generation strategy to generate diverse 3D shapes.
+After pretraining, following Li et al.
+(2023), we employ an iterative generation strategy to generate diverse 3D shapes.
 The generation process starts with all tokens being considered as masked tokens.
@@
 Specifically, in each iteration, based on the predicted probabilities from the generator, we sample and replace those tokens with higher probabilities.
-Table 1: Shape classification results on ModelNet40 Wu_2015_CVPR and ScanObjectNN Uy_2019_ICCV .
-We report the accuracy (%) of both training from scratch and fine-tuning models on ModelNet40 and ScanObjectNN datasets.
-Methods Reference ModelNet40 ScanObjectNN
-OBJ_BG OBJ_ONLY PB_T50_RS
-training from scratch
-PointNet Qi_2017_CVPR CVPR 2017 89.2 73.3 79.2 68.0
-PointNet++ qi2017pointnet++ Neurips 2017 90.5 82.3 84.3 77.9
-PointCNN li2018pointcnn Neurips 2018 92.2 86.1 85.5 78.5
-DGCNN wang2019dynamic TOG 2019 92.9 82.8 86.2 78.1
-GBNet 9410405 TMM 2021 93.8 - - -
-MVTN Hamdi_2021_ICCV ICCV 2021 93.8 92.6 - 82.8
-PointNeXt qian2022pointnext Neurips 2022 94.0 - - -
-fine-tuning
-Point-BERT Yu_2022_CVPR CVPR 2022 93.2
-Mask-Point liu2022masked ECCV 2022 93.8 89.3 88.1 84.3
-Point-MAE pang2022masked ECCV 2022 93.8 90.0 88.2 85.2
-Point-M2AE zhang2022point Neurips 2022 94.0 91.2 88.8 86.4
-PointGPT NEURIPS2023_5ed5c3c8 Neurips 2023 94.0 91.6 90.0 86.9
-Point-MAGE - 94.2 92.9 91.0 87.1
-Table 2: Few-shot classification on ModelNet40Wu_2015_CVPR . We report the average accuracy (%) and standard deviation (%) of 10 independent experiments.
-Methods Reference 5-way 10-way
+4 Experiments
+In this section, we first provide a detailed overview of the pretraining setup of Point-MGE.
+We then evaluate the pretrained model’s performance in both representation learning and generation capabilities.
+Finally, to gain a deeper understanding of the roles and importance of each component in the model, we conduct ablation studies on Point-MGE.
+More results can be found in the Appendix.
+Methods 5-way 10-way
 10-shot 20-shot 10-shot 20-shot
-training from scratch
-DGCNN wang2019dynamic TOG 2019 31.6 ± 2.8 40.8 ± 4.6 19.9 ± 2.1 16.9 ± 1.5
-OcCo Wang_2021_ICCV ICCV 2021 90.6 ± 2.8 92.5 ± 1.9 82.9 ± 1.3 86.5 ± 2.2
-fine-tuning
-Transformer + OcCo Yu_2022_CVPR CVPR 2022 94.0 ± 3.6 95.9 ± 2.3 89.4 ± 5.1 92.4 ± 4.6
-Point-BERT Yu_2022_CVPR CVPR 2022 94.6 ± 3.1 96.3 ± 2.7 91.0 ± 5.4 92.7 ± 5.1
-Mask-Point liu2022masked ECCV 2022 95.0 ± 3.7 97.2 ± 1.7 91.4 ± 4.0 93.4 ± 3.5
-Point-MAE pang2022masked ECCV 2022 96.3 ± 2.5 97.8 ± 1.8 92.6 ± 4.1 95.0 ± 3.0
-Point-M2AE zhang2022point Neurips 2022 96.8 ± 1.8 98.3 ± 1.4 92.3 ± 4.5 95.0 ± 3.0
-PointGPT NEURIPS2023_5ed5c3c8 Neurips 2023 96.8 ± 2.0 98.6 ± 1.1 92.6 ± 4.6 95.2 ± 3.4
-Point-MAGE - 97.2 ± 1.9 98.6 ± 0.7 92.7 ± 4.0 95.4 ± 3.0
-4 Experiments
-In this section, we first provide a detailed overview of the pretraining setup of Point-MAGE.
-We then evaluate the pretrained model’s performance in both representation learning and generation capabilities.
-Finally, to gain a deeper understanding of the roles and importance of each component in the model, we conduct ablation studies on Point-MAGE.
-More results can be found in the Appendix.
+Supervised Learning Only
+PointNet 52.0±3.8 57.8±4.9 46.6±4.3 35.2±4.8
+PointNet-OcCo 89.7±1.9 92.4±1.6 83.9±1.8 89.7±1.5
+DGCNN 31.6±2.8 40.8±4.6 19.9±2.1 16.9±1.5
+with Single-Modal Pre-Training
+Point-BERT 94.6±3.1 96.3±2.7 91.0±5.4 92.7±5.1
+Mask-Point 95.0±3.7 97.2±1.7 91.4±4.0 93.4±3.5
+Point-MAE 96.3±2.5 97.8±1.8 92.6±4.1 95.0±3.0
+Point-M2AE 96.8±1.8 98.3±1.4 92.3±4.5 95.0±3.0
+PointGPT 96.8±2.0 98.6±1.1 92.6±4.6 95.2±3.4
+GPM 97.2±2.6 98.7±2.2 92.9±4.2 95.0±3.0
+Point-MGE 97.2±1.7 98.7±0.7 92.8±4.0 95.4±3.2
+Improvement +2.6 +2.4 +1.8 +2.7
+PointGPT-B 97.5±2.0 98.8±1.0 93.5±4.0 95.8±3.0
+Point-MGE-B 97.7±1.8 98.8±0.8 93.8±4.6 95.8±3.4
+with Cross-Modal Pre-Training
+ACT 96.8±2.3 98.0±1.4 93.3±4.0 95.6±2.8
+Joint-MAE 96.7±2.2 97.9±1.8 92.6±3.7 95.1±2.6
+I2P-MAE 97.0±1.8 98.3±1.3 92.6±5.0 95.5±3.0
+Recon 97.3±1.9 98.9±1.2 93.3±3.9 95.8±3.0
+Table 2: Few-shot classification on ModelNet40 Wu et al. (2015). We report the average accuracy (%) and standard deviation (%) of 10 independent experiments.
 4.1 Pre-training Setups
-Point-MAGE is pre-trained on the ShapeNet chang2015shapenet dataset, which includes 55 different object categories and over 50,000 3D models.
-From the surface of each 3D shape, 2048 points are uniformly sampled and further divided into 512 patches, each containing 32 points.
-In the first stage, we pre-trained the VQVAE to obtains a discrete token for each point patch, resulting in a token sequence of 512 tokens.
-In the second stage, we determine the variable masking ratio with $\mu=0.55$ and $\sigma=0.25$ and pre-train the extractor-generator architecture based on masked token reconstruction.
-Additionally, to validate the capability of integrating the Point-MAGE framework with advanced model designs, we pre-trained a multi-scale extractor model following zhang2022point and evaluated its representation learning performance across various downstream tasks.
-We used the AdamW optimizer for pre-training the two stages for 800 and 300 epochs, respectively.
+Point-MGE is pre-trained on the ShapeNet Chang et al.
+(2015) dataset, which includes 55 different object categories and over 50,000 3D models.
+From the surface of each 3D shape, 2048 points are uniformly sampled and further divided into 64 patches, each containing 32 points.
+First, we trained the VQVAE to obtains the discrete tokens.
+Next, we determined the sliding masking ratio, setting $\beta=0.5$ and $u=2$, and pre-trained the ViT-based extractor-generator architecture using multi-scale masked token reconstruction.
+The number of scales and the number of point patches at each scale were consistent with Point-M2AE Zhang et al.
+(2022b).
+Additionally, to validate the high-capacity model learning ability of the Point-MGE framework, we scaled the model to ViT-B as described in Chen et al.
+(2023), performed post-training, and evaluated its representation learning performance across various downstream tasks.
+We used the AdamW optimizer to perform pre-training for 300 epochs.
 Further implementation details and hyperparameter settings are provided in the Appendix.
@@
 To demonstrate the representation learning performance of our method, we conducted experiments on shape classification, few-shot learning, and part segmentation.
-Additionally, the Appendix will provide further research and discussion on the properties of Point-MAGE.
+We evaluated the benchmark performance for supervised learning only, single-modal pre-training, and cross-modal pre-training separately.
 Shape Classification.
 We conducted shape classification fine-tuning on two datasets.
-The ModelNet40 Wu_2015_CVPR dataset includes 3D CAD models of 40 categories, and we followed the official data split provided in  Wu_2015_CVPR for our experiments.
-As shown in Table 1, Point-MAGE achieved a classification accuracy of 94.2% on this dataset, surpassing other existing pre-trained models.
-The ScanObjectNN  Uy_2019_ICCV dataset, derived from real-world indoor scans, contains approximately 15,000 objects across 15 categories.
+The ModelNet40 Wu et al.
+(2015) dataset includes 3D CAD models of 40 categories, and we followed the official data split provided in  Wu et al.
+(2015) for our experiments.
+As shown in Table 1, Point-MGE achieved a second-place classification accuracy of 94.2% on this dataset.
+The ScanObjectNN  Uy et al.
+(2019)dataset, derived from real-world indoor scans, contains approximately 15,000 objects across 15 categories.
 We evaluated this dataset using three different data splits: OBJ-BG, OBJ-ONLY, and PB-T50-RS.
-The experimental results indicate that Point-MAGE consistently outperformed other models across all splits.
-Notably, in the OBJ-BG setting, Point-MAGE achieved an accuracy of 92.9%, outperforming the second-best pre-trained model NEURIPS2023_5ed5c3c8 by 1.3 percentage points.
+The experimental results indicate that Point-MGE consistently outperformed other models across all splits.
+Notably, in the OBJ-BG setting, Point-MGE achieved an accuracy of 92.9%, outperforming the second-best pre-trained model Chen et al.
+(2023) by 1.3 percentage points.
+Additionally, we extended Point-MGE and three benchmark methods to the ViT-B model for performance evaluation.
+The results demonstrate that under high-capacity model conditions, our method still maintains a competitive advantage.
+Methods $\text{mIoU}_{C}$uparrow $\text{mIoU}_{I}$uparrow
+Supervised Learning Only
+PointNet Qi et al. (2017a) 80.4 83.7
+PointNet++ Qi et al. (2017b) 81.9 85.1
+DGCNN Wang et al. (2019) 82.3 85.2
+with Single-Modal Pre-Training
+Transformer + OcCo Yu et al. (2022) 83.4 85.1
+Point-BERT Yu et al. (2022) 84.1 85.6
+Mask-Point Liu et al. (2022) 84.4 86.0
+Point-MAE Pang et al. (2022) - 86.1
+PointGPT Chen et al. (2023) 84.1 86.2
+GPM Li et al. (2024) 84.2 85.8
+Point-MGE 85.0 86.5
+Improvement +0.9 +0.9
+PointGPT-B Chen et al. (2023) 84.5 86.5
+Point-MGE-B 85.0 86.6
+with Cross-Modal Pre-Training
+ACT Dong et al. (2023) 84.7 86.1
+Recon Qi et al. (2023) 84.8 86.4
+Table 3: Part segmentation on ShapeNetPart Yi et al.
+(2016).
+We report the mean IoU (%) of all part categories (’$\text{mIoU}_{C}$’) and the mean IoU of all instances (’$\text{mIoU}_{I}$’) in the dataset.
 Few-shot Classification.
-We evaluated the few-shot learning performance of Point-MAGE by conducting N-way, K-shot experiments on the ModelNet40Wu_2015_CVPR dataset.
+We evaluated the few-shot learning performance of Point-MGE by conducting N-way, K-shot experiments on the ModelNet40Wu et al.
+(2015) dataset.
 Specifically, we randomly selected N classes from ModelNet40 and sampled K objects from each class.
@@
 We evaluated each of the four combinations of settings with N = {5, 10} and K = {10, 20}.
-As shown in the table 2, we achieved SOTA performance in all of them, demonstrating that Point-MAGE’s pre-trained 3D representation is well-suited for few-shot learning tasks.
-Table 3: Part segmentation on ShapeNetPart yi2016scalable .
-We report the mean IoU (%) of all part categories (’$\text{mIoU}_{C}$’) and the mean IoU of all instances (’$\text{mIoU}_{I}$’) in the dataset.
-       Methods        Reference        $\text{mIoU}_{C}$uparrow        $\text{mIoU}_{I}$uparrow
-       training from scratch
-       PointNet Qi_2017_CVPR        CVPR 2017        80.4        83.7
-       PointNet++ qi2017pointnet++        Neurips 2017        81.9        85.1
-       DGCNN wang2019dynamic        TOG 2019        82.3        85.2
-       fine-tuning
-       Transformer + OcCo Yu_2022_CVPR        CVPR 2022        83.4        85.1
-       Point-BERT Yu_2022_CVPR        CVPR 2022        84.1        85.6
-       Mask-Point liu2022masked        ECCV 2022        84.4        86.0
-       Point-MAE pang2022masked        ECCV 2022        -        86.1
-       Point-M2AE zhang2022point        Neurips 2022        84.9        86.5
-       PointGPT NEURIPS2023_5ed5c3c8        Neurips 2023        84.1        86.2
-       Point-MAGE        -        85.0        86.5
+As shown in Table 2, under both ViT and ViT-B settings, we achieved SOTA performance in at least three combinations, demonstrating that Point-MGE’s pre-trained 3D representations are highly effective for few-shot learning tasks.
+Part Segmentation.
+Table 3 presents the results of average IoU for all categories ($\text{mIoU}_{C}$) and all instances ($\text{mIoU}_{I}$) in the part segmentation experiments on the ShapeNetPart Yi et al.
+(2016) dataset.
+Compared to the fully supervised method DGCNNWang et al.
+(2019), Point-MGE achieved significant improvements of 2.7% and 1.3% in these two key evaluation metrics, respectively.
+Our method also outperforms other pretrained models, achieving a $\text{mIoU}_{C}$ of 85.0%, which ranks first.
+Additionally, the experimental results show that increasing the number of parameters yields only marginal improvements for this task.
+4.3 3D Shape Generation
+Methods S-FPD $\downarrow$ S-KPD $\downarrow$ R-FID $\downarrow$ R-KID $\downarrow$
+(×$\times 10^{3}$) (×$\times 10^{3}$)
+Grid-$8^{3}$ 4.03 6.15 32.78 14.12
+PVD 2.33 2.65 270.64 281.5
+3DILG 1.89 2.17 24.83 10.51
+3DShape2VecSet 0.97 1.11 24.24 11.76
+Point-MGE 0.85 1.34 21.63 11.48
 Table 4: Unconditional generation. We report four metrics for evaluating the unconditional shape generation task.
-Methods Surface-FPD $\downarrow$ Surface-KPD (×$10^{3}$$\downarrow$ Rendering-FID  $\downarrow$ Rendering-KID (×$10^{3}$$\downarrow$
-Grid-$8^{3}$ zhang20223dilg 4.03 6.15 32.78 14.12
-3DILG zhang20223dilg 1.89 2.17 24.83 10.51
-Point-MAGE 0.98 1.34 22.79 11.48
-Figure 2: Visualization of unconditional generation. All models were trained on the complete ShapeNet dataset, and we sampled nine generated results for rendering for each model.
-Part Segmentation.
-Table 3 presents the results of average IoU for all categories ($\text{mIoU}{C}$) and all instances ($\text{mIoU}{I}$) in the part segmentation experiments on the ShapeNetPart yi2016scalable dataset.
-Compared to the fully supervised method DGCNNwang2019dynamic , Point-MAGE achieved significant improvements of 2.7% and 1.3% in these two key evaluation metrics, respectively.
-Our method also outperforms other pretrained models, achieving a $\text{mIoU}_{C}$ of 85.0%, which ranks first.
-4.3 3D Shape Generation
+Figure 2: Visualization of unconditional generation. All models were trained on the complete ShapeNet dataset, and we sampled nine generated results for rendering.
 Unconditional Generation.
-After completing the pre-training stage, Point-MAGE is capable of performing unconditional generation tasks.
+After completing the pre-training stage, Point-MGE is capable of performing unconditional generation tasks.
 fréchet inception distance (FID) and kernel inception distance (KID) are commonly used to evaluate the generation quality of generative models.
-Following zhang20233dshape2vecset , we use Rendering-FID and Rendering-KID metrics to evaluate the FID and KID of rendered images of 3D shapes, as well as Surface-FPD and Surface-KPD metrics to evaluate the FID and KID of sampled point clouds from 3D shapes.
+Following Zhang et al.
+(2023a), we use Rendering-FID (R-FID) and Rendering-KID (R-KID) metrics to evaluate the FID and KID of rendered images of 3D shapes, as well as Surface-FPD (S-FPD) and Surface-KPD (S-KPD) metrics to evaluate the FID and KID of sampled point clouds from 3D shapes.
 Table 4 compares the performance of various methods in unconditional shape generation on ShapeNet.
-3DILG also adopts a two-stage generation strategy and completes the generation training through an autoregressive process.
-Point-MAGE outperforms other methods on all three evaluation metrics, indicating that our method is capable of generating 3D shapes that match the real distribution.
-Additionally, we visualize the rendered results of generated shapes for all methods in Figure 2.
+3DILG Zhang et al.
+(2022a) also adopts a two-stage generation strategy and completes the generation training through an autoregressive process.
+PVD Zhou et al.
+(2021) and 3DShape2VecSet Zhang et al.
+(2023a) are generative methods based on diffusion models.
+Point-MGE outperforms other methods on two of the evaluation metrics, indicating that our approach can generate 3D shapes that closely match the real distribution.
+Additionally, in Figure  2, we visualize the rendered results of the generated shapes.
+Methods Airplane Faucet Table Chair
+S-FID $\downarrow$ 3DILG 0.71 2.49 2.10 0.96
+NeuralWavelet 0.38 1.84 1.12 1.14
+3DShape2VecSet 0.62 1.47 1.19 0.76
+Point-MGE 0.67 1.65 1.03 0.71
+S-KID $\downarrow$ (×$\times 10^{3}$) 3DILG 0.81 3.76 3.84 1.21
+NeuralWavelet 0.53 2.41 1.55 1.50
+3DShape2VecSet 0.83 1.34 1.87 0.70
+Point-MGE 0.79 1.31 1.14 0.58
+Table 5: Category-conditional generation. We reported the S-FID and S-KID metrics for four categories.
 Figure 3: Visualization of category-conditional generation. We selected three category labels (airplane, faucet, and table) to display the rendered results of the generated shapes.
@@
 The generated results demonstrated the ability of different models to reconstruct the complete shapes.
+Extractor Generator Acc (%) R-FID $\downarrow$
+6 6 91.2 29.07
+8 6 91.9 27.85
+12 8 92.4 21.63
+12 12 92.1 20.94
+Table 6: Effects of varying transformer depths. The accuracy (%) on the ScanObjectNN Uy et al. (2019) (OBJ_BG) and the R-FID for unconditional generation are reported.
 Conditional Generation.
-We evaluated the conditional generation capabilities of Point-MAGE under two different settings.
-First, we retrained the model using the 54 categories from the ShapeNet dataset as conditional embeddings.
+We evaluated the conditional generation capabilities of Point-MGE under two different settings.
+First, we retrained the model using the 54 categories from the ShapeNet Chang et al.
+(2015) dataset as conditional embeddings.
+As shown in Table 5, we evaluated the Surface-FID (S-FID) and Surface-KID (S-KID) for our method and the benchmark methods across four categories.
+The results indicate that our method achieved performance comparable to or even surpassing those of purely generative approaches.
 Figure 3 displays the rendered results of generated shapes for three categories (airplane, faucet, and table), with nine instances shown for each category.
-These visualizations demonstrate our method’s ability to generate realistic 3D shapes based on categorical conditions.
-Next, we further evaluated Point-MAGE against other SSL models in the masked generation task.
+Next, we further evaluated Point-MGE against other SSL models in the masked generation task.
 As shown in Figure 4, these models typically train by reconstructing local point clouds, which primarily allows them to complete point cloud reconstruction tasks and depends on the provision of the central point positions.
@@
 In contrast, our method can reconstruct 3D shape meshes using the minimal amount of point cloud information and does not rely on masked center point information during the generation process.
+Type Acc (%) R-FID $\downarrow$
+Quantized features 92.4 21.63
+Index-embedded tokens 87.2 20.34
+Table 7: Effects of input type. The accuracy (%) and the R-FID are reported.
+Sliding Masking Various Masking
+$\beta=0.5$ $\beta=0.5$ $\beta=0.7$ $\mu=1.0$ $\mu=0.7$
+$u=1$ $u=2$ $u=2$ $\sigma=0$ $\sigma=0.25$
+Acc (%) 91.9 92.4 91.7 86.2 91.6
+R-FID $\downarrow$ 24.25 21.63 22.71 26.72 22.43
+Table 8: Effects of masking design.The accuracy (%) and the R-FID are reported.
+NeRFs EMD CD L1 CD L2
+Acc (%) 92.4 87.9 88.7 89.1
+R-FID $\downarrow$ 21.63 N/A N/A N/A
+Table 9: Effects of input type. The accuracy (%) and the R-FID are reported.
 4.4 Ablation Studies
-In this section, we discuss three aspects that influence the representation learning and generation capabilities of the pre-trained model: the transformer design of the extractor-generator architecture, the impact of different inputs, and the effect of varying masking ratios.
+In this section, we discuss several factors that influence the representation learning and generative capabilities of the pre-trained model.
 Transformer Depth.
-As shown in Table 5(a), we used the standard ViT architecture to build the extractor-generator structure and explored four different transformer depth combinations.
+As shown in Table 6, we used the standard single scale ViT architecture to build the extractor-generator structure and explored four different transformer depth combinations.
 The results indicate that increasing the model depth significantly enhances the performance of the pre-trained model.
-Additionally, consistent with trends observed in previous masked modeling studies he2022masked ; pang2022masked ; zhang2022point , an asymmetric extractor-generator structure is more beneficial for representation learning.
+Additionally, consistent with trends observed in previous masked modeling studies He et al.
+(2022); Pang et al.
+(2022); Zhang et al.
+(2022b), an asymmetric extractor-generator structure is more beneficial for representation learning.
 This reliance on the extractor to obtain high-quality representations from a small number of visible tokens is due to the use of a relatively simple generator to complete the reconstruction task.
 Input Type.
-In Table 5(b), we investigate the impact of two different input types on model performance.
+In Table 7, we investigate the impact of two different input types on model performance.
 Specifically, after vector quantizing the initial features of the point cloud patches, we consider using either the quantized features or their indices in the codebook as inputs.
 Experimental results show that using index-embedded tokens benefits the generation task, but its generalization ability across datasets is limited, which is detrimental to fine-tuning on downstream tasks, leading to poorer representation learning capabilities.
-This contrasts sharply with the results shown by MAGE Li_2023_CVPR , where using index embeddings performs well in both aspects.
-We believe this characteristic of MAGE arises from its pre-training on a larger-scale dataset, and the advantages of image data over point cloud data in terms of regularity and consistency, which reduce the impact of different data sources.
 Masking Design.
-To determine the optimal masking ratio, in Table 5(c), we compare the impact of various masking ratio distributions on fine-tuning and near-unconditional generation task results.
-Here, $\mu$ and $\sigma$ represent the mean and standard deviation of the truncated Gaussian distribution, respectively.
-When $\sigma=0$, a fixed masking ratio is used.
-In this case, when $\mu=1.0$, all tokens are masked, preventing the extractor from obtaining useful information, making the fine-tuning process essentially equivalent to training from scratch.
-In other cases, since the generator does not experience the process of reconstructing from zero, it fails to complete the generation task, and the generation results are marked as not applicable (N/A).
-The findings indicate that employing a variable masking ratio not only enables the completion of generation tasks but also encourages the representation learning process to capture superior features.
-Table 5: Ablation experiments. The fine-tuned accuracy on the ScanObjectNN Uy_2019_ICCV (OBJ_BG) and the Rendering-FID (R-FID) for unconditional generation are reported.
-(a) Transformer depth
-Extractor Generator Acc (%) R-FID $\downarrow$
-6 6 91.1 29.07
-8 6 91.7 27.85
-12 8 92.4 22.79
-12 12 91.9 20.94
-(b) Input type
-Type Acc (%) R-FID $\downarrow$
-Quantized features 92.4 22.79
-Index-embedded tokens 87.3 22.36
-(c) Masking ratio distribution
-$\mu=1.0$ $\mu=0.7$ $\mu=0.6$ $\mu=0.55$ $\mu=0.5$ $\mu=0.55$ $\mu=0.55$ $\mu=0.55$
-$\sigma=0$ $\sigma=0$ $\sigma=0$ $\sigma=0$ $\sigma=0$ $\sigma=0.15$ $\sigma=0.25$ $\sigma=0.5$
-Acc (%) 86.2 91.8 91.9 92.0 91.8 92.2 92.4 92.1
-R-FID $\downarrow$ 26.72 N/A N/A N/A N/A 23.58 22.79 23.65
+In determining the optimal masking strategy, as shown in Table 8, we compared our sliding masking strategy with the variable masking strategy used in MAGE Li et al.
+(2023) and analyzed the impact of different hyperparameters on performance.
+For the variable masking, $\mu$ and $\sigma$ represent the mean and standard deviation of the truncated Gaussian distribution, respectively.
+The results indicate that the variable masking ratio, which performs well on 2D images, does not provide an advantage in the point cloud modality, where data is relatively scarce.
+In contrast, our smooth masking strategy better facilitates the transition between representation learning and generative training.
+Reconstruction Target.
+In exploring the training of VQVAE, we compared different reconstruction targets.
+As shown in Table 9, this study employs NeRFs as the reconstruction target.
+Additionally, we experimented with directly using point clouds as the reconstruction target and tested methods using Earth Mover’s Distance (EMD) and Chamfer Distance (CD) as loss functions.
+The experimental results demonstrate that our approach excels in overcoming point cloud sampling bias, leading to superior representation learning capabilities.
+However, directly using point clouds as the reconstruction target, similar to previous methods, is limited to generating point clouds and cannot produce high-resolution 3D shapes.
+In the table, this generative capability is marked as Not Applicable (N/A).
 5 Conclusion
-We propose Point-MAGE, a novel unified framework for shape generation and representation learning on point clouds.
-Our approach serializes point cloud segments and employs a two-stage process to obtain quantized tokens, applying a variable masking ratio to accommodate both tasks simultaneously.
-We have achieved extensive results in shape classification, few-shot transfer learning, unconditional shape generation, and conditional generation.
-In terms of representation learning, our method has set a new SOTA result.
-In the future, we plan to extend Point-MAGE to pretrain on larger datasets to achieve even better performance.
-We hope our work draws the attention of researchers to the synergistic effects of combining two types of models in 3D vision.
-Appendix
-Appendix A Implementation Details
-Our pre-trained model was trained on four A800 80G GPUs, and the fine-tuned model was trained on two 3090 24G GPUs.
-Below, we provide detailed descriptions of the training procedures and hyperparameter settings for each task.
-Positional Encodings.
-For absolute position encoding, we use a two-layer multilayer perceptron (MLP) to encode the center point of each point cloud patch into a D-dimensional vector.
-For relative position encoding, we use a learnable embedding layer to embed the indices of the fixed-length point patch sequence.
-Pre-training.
-In the ShapeNet [4] dataset, we randomly sampled each 3D shape to generate a point set of 2048 points, which served as input data for the two-stage model training.
-To improve the model’s generalization ability, we employed data augmentation techniques, including random scaling and flipping of the sampled point clouds.
-In the first stage, the point set was divided into 512 point patches, each containing 32 points.
-The VQVAE’s encoder and decoder have a ViT model depth of 6.
-The training parameters were set as follows: a batch size of 128, an initial learning rate of $5\times 10^{-4}$, and a weight decay coefficient of $5\times 10^{-2}$.
-In the second stage, the point patches were tokenized to obtain a token sequence with a length of 512.
-The extractor’s depth was set to 12, and the generator’s depth was set to 8.
-Additionally, we pre-trained a multi-scale extractor with a scale value of 3, where the number of point patches for each scale was [512, 256, 64], and the corresponding token dimensions were [96, 192, 384].
-The training parameters for this stage were set as follows: a batch size of 256, an initial learning rate of $10^{-3}$, and a weight decay coefficient of $5\times 10^{-2}$.
-In both stages, we used a cosine annealing learning rate scheduler and employed a warm-up phase during the first 10 epochs of training.
-Shape Classification.
-For the shape classification task, we finetuned the Point-MAGE model on two datasets: ModelNet40[45] and ScanObjectNN[36].
-The ModelNet40[45] dataset consists of synthetic 3D shapes from 40 categories.
-We used 9843 samples for training and 2468 samples for validation.
-For each sample, we sampled 2048 points to construct the input for the model.
-On the other hand, the ScanObjectNN[36] dataset contains real-world 3D shapes from 15 categories, divided into three subsets: OBJ-BG, OBJ-ONLY, and PB-T50-RS, with a total of 11416 training point clouds and 2882 validation point clouds.
-For these point cloud samples, we sampled 2048 points as input.
-During the finetuning training, both datasets were trained for 300 epochs, with a batch size of 32, learning rate of $5\times 10^{-4}$, and weight decay coefficient of $5\times 10^{-2}$.
-We kept the other training hyperparameters consistent with the settings from the pretraining stage.
-Few-shot Classification.
-We randomly selected K categories from the 40 categories in the ModelNet40[45] dataset and sampled N + 20 3D shapes for each selected category, where N shapes were used for model training and the remaining 20 shapes were used for testing.
-We designed four different few-shot learning settings: 5-way 10-shot, 5-way 20-shot, 10-way 10-shot, and 10-way 20-shot.
-Except for the training epochs set to 150, other training settings remained consistent with the shape classification experiments.
-Part Segmentation.
-We conducted experiments on the ShapeNetPart[53] dataset, which includes 16 object categories and 50 part categories, totaling 16,881 3D shape models.
-We selected 14,007 shapes from the dataset for training and reserved 2,874 shapes for validation.
-During training, the model was optimized for 300 epochs, with each batch processing 16 input point clouds, each containing 2048 points.
-The learning rate was set to $2\times 10^{-4}$, and the weight decay coefficient was 0.1.
-Appendix B Additional Results and Analysis
-Learning Curves.
-As shown in Figure 5, we compared the learning curves of our method and Point-M2AE [57] during the pre-training phase for reconstructing each local point cloud patch.
-We used the same single-layer convolutional reconstruction head as Point-M2AE for reconstructing each point cloud patch.
-It is noteworthy that our pre-training process optimizes the token reconstruction loss rather than the traditional point cloud Chamfer distance loss.
-Surprisingly, even without using the corresponding loss function, our method can still effectively learn to reconstruct local point cloud patches by reconstructing semantic tokens.
-This finding reveals that the feature information learned through point cloud reconstruction is actually a subset of the feature information learned by our method.
-Ultimately, we showed the difference in loss values near convergence: our method’s reconstruction loss is slightly higher than that of Point-M2AE.
-We believe this difference is due to the point cloud sampling variance issue discussed in the main text.
-Using the Chamfer distance loss forces the model to learn specific surface sampling, which may lead to a lower loss value upon convergence but does not benefit the model in learning the general characteristics of surface shapes.
-Figure 5: Visualization of masked point clouds conditional generation.
-Performance of Linear SVM.
-As shown in Table 6, we provide detailed reports on the linear SVM classification accuracy of Point-MAGE and other SSL methods on the ModelNet40[45] dataset, and find that our method achieves outstanding performance.
-Specifically, compared to the pioneering method Point-BERT[54] for point cloud masking SSL, our model achieves a significant improvement of 5.7% in classification accuracy.
-Table 6: Linear SVM Performance on ModelNet40[45]. We report the shape classification accuracy (%).
-Methods Reference Acc (%)
-3D-GAN[44] Neurips 2016 83.3
-Latent-GAN[38] TMM 2020 85.7
-FoldingNet[52] CVPR 2018 88.4
-MAP-VAE[14] ICCV 2019 88.4
-VIP-GAN[13] AAAI 2019 90.2
-DGCNN + Jiasaw[33] Neurips 2019 90.6
-DGCNN + OcCo[41] ICCV 2021 90.7
-DGCNN + CrossPoint[2] CVPR 2022 91.2
-Transformer + OcCo[54] CVPR 2022 89.6
-Point-BERT[54] CVPR 2022 87.4
-Point-MAE[25] ECCV 2022 91.0
-Point-M2AE[57] Neurips 2022 92.9
-Point-MAGE - 93.1
-t-SNE Visualizations.
-In Figure 6, we applied t-SNE to visualize the feature distributions on the ModelNet40 [45] dataset, showing the results of three different training strategies: (a) training from scratch, (b) pre-training, and (c) fine-tuning.
-The observations reveal that when the Point-MAGE model is trained from scratch, there is significant overlap of features from different categories in the feature space.
-However, after pre-training, the model can generate more distinguishable features for different categories, enhancing feature recognizability.
-Furthermore, with fine-tuning, we found that the features of samples within the same category become more tightly clustered in the feature space, while the features of samples from different categories show clear separation.
-This indicates that the fine-tuning process further optimizes the clustering of features and enhances the model’s ability to distinguish between features of different categories.
-Figure 6: t-SNE visualization on ModelNet40[45].
-Part Segmentation.
-In the part segmentation task on the ShapeNetPart dataset [53], the Point-MAGE model demonstrated outstanding performance.
-As shown in Table 7, Point-MAGE achieved the highest accuracy in 10 out of 16 categories.
-Overall, the average accuracy of Point-MAGE was the best among all methods, proving its effectiveness in handling complex 3D shapes and part segmentation tasks.
-Table 7: Additional results of part segmentation on ShapeNetPart[53]. We report the IoU (%) for each category.
-Methods aero bag cap car chair earp. guit. knif. lamp lapt. mot. mug pist. rock. skt. table
-PointNet[26] 83.4 78.7 82.5 74.9 89.6 73.0 91.5 85.9 80.8 95.3 65.2 93.0 81.2 57.9 72.8 80.6
-PointNet++[27] 82.4 79.0 87.7 77.3 90.8 71.8 91.0 85.9 83.7 95.3 71.6 94.1 81.3 58.7 76.4 82.6
-DGCNN[42] 84.0 83.4 86.7 77.8 90.6 74.7 91.2 87.5 82.8 95.7 66.3 94.9 81.1 63.5 74.5 82.6
-Point-BERT[54] 84.3 84.8 88.0 79.8 91.0 81.7 91.6 87.9 85.2 95.6 75.6 94.7 84.3 63.4 76.3 81.5
-Mask-Point[21] 84.2 85.6 88.1 80.3 91.2 79.5 91.9 87.8 86.2 95.3 76.9 95.0 85.3 64.4 76.9 81.8
-Point-MAE[25] 84.3 85.0 88.3 80.5 91.3 78.5 92.1 87.4 86.1 96.1 75.2 94.6 84.7 63.5 77.1 82.4
-Point-MAGE 85.2 87.2 89.2 81.4 91.8 80.4 92.4 88.4 86.0 96.0 78.0 95.3 85.0 64.1 77.6 82.0
-Appendix C Limitations and Future Directions
-Despite the significant enhancements our method brought to the capabilities of the pre-trained model, these improvements did not translate into substantial performance gains across all downstream tasks.
-For instance, in the part segmentation experiments, the observed performance gain was relatively small.
-Additionally, the performance of our method was influenced by the input type (quantized features or index embeddings), and it did not fully replicate the success of MAGE [19] in achieving optimal representation learning and generative performance simultaneously in the 2D image domain.
-We attribute this phenomenon partly to the fact that the scale of the point cloud pre-training dataset is much smaller than that of the 2D image domain.
-Therefore, we expect that by expanding the pre-training dataset, we can further improve model performance and achieve a better unification of representation learning and generative performance.
+We present Point-MGE, a novel unified framework for point cloud shape generation and representation learning.
+Our approach serializes point patches to obtain quantized tokens and applies a sliding masking ratio, allowing it to adapt to both representation learning and generation tasks.
+Our method excels in addressing the sampling bias issue in point clouds, setting new SOTA results in shape classification, few-shot learning, and part segmentation tasks.
+Additionally, in generation tasks, our approach achieves near SOTA performance in both unconditional and conditional shape generation, producing realistic high-resolution 3D shapes.
+We hope our work will draw attention from researchers to the synergistic effects of combining these two model types in 3D vision.
+Acknowledgments
+This work was supported by the Guangdong Major Project of Basic and Applied Basic Research (2023B0303000016).